### v1

# RocketKV: Accelerating Long-Context LLM Inference via Two-Stage KV Cache Compression

Abstract
Transformer-based Large Language Models rely critically on KV cache to efficiently handle extended contexts during the decode phase. Yet, the size of the KV cache grows proportionally with the input length, burdening both memory bandwidth and capacity as decoding progresses. To address this challenge, we present RocketKV, a training-free KV cache compression strategy designed specifically to reduce both memory bandwidth and capacity demand of KV cache during the decode phase. RocketKV contains two consecutive stages. In the first stage, it performs coarse-grain KV cache eviction on the input sequence tokens with SnapKV++, a method improved upon SnapKV by introducing adaptive pooling size and full compatibility with grouped-query attention. In the second stage, it adopts a hybrid attention method to conduct fine-grain top-k sparse attention, approximating the attention scores by leveraging both head and sequence dimensional reductions. Combining these two stages, RocketKV achieves significant KV cache fetching bandwidth and storage savings while maintaining comparable accuracy to full KV cache attention. We show that RocketKV provides end-to-end speedup by up to 3$\times$ as well as peak memory reduction by up to 31% in the decode phase on an NVIDIA H100 GPU compared to the full KV cache baseline, while achieving negligible accuracy loss on a variety of long-context tasks.
Payman Behnam* $\circ$ 12 Yaosheng Fu*1 Ritchie Zhao1 Po-An Tsai1 Zhiding Yu1 Alexey Tumanov2
footnotetext: Corresponding authors: Payman Behnam, payman.behnam@gatech.edu.edu; Yaosheng Fu, yfu@nvidia.com
*: Equal contribution
$\circ$: Work done during Payman’s internship at NVIDIA Research
1NVIDIA, Santa Clara, USA
2Georgia Institute of Technology, Atlanta, USA
1 Introduction
Long-context Large Language Models (LLMs) have enhanced the capabilities of LLMs across diverse tasks that require processing extensive amounts of text. Examples of such applications are document summarization, legal document analysis, coding assistants, conversational AI, etc [24, 22, 12].
In transformer-based LLM [33] inference, the key-value cache (KV cache)—which stores past attention keys and values to avoid recomputation—becomes a major bottleneck during the decode phase, as its size scales linearly with both the sequence length and batch size. For example, the Llama3.1-70B-Instruct [1] model with a batch size of 32, and a context length of 32K requires around 320GB of KV cache storage at FP16 precision, which even advanced hardware [23, 15, 3] can hardly handle.
Fortunately, previous work [37, 31, 25, 20, 28, 9] has shown that only a small subset of KV tokens is required at each decode step to maintain accuracy. Therefore, if those KV tokens can be accurately predicted in advance, dense attention operations can be replaced with sparse attention operations with significant memory bandwidth and capacity improvement. These methods often fall into two categories: 1) permanent KV token eviction, and 2) dynamic KV token selection. The former results in both memory bandwidth and storage savings, but could lead to noticeable accuracy loss if KV tokens dropped earlier are needed by later decode steps. The latter avoids this shortcoming by keeping all KV tokens in the memory and dynamically selecting a subset each time. Hence, it only results in memory bandwidth savings, but often requires extra memory storage overhead for auxiliary data.
Aiming to achieve better trade-offs among memory bandwidth, capacity, and model accuracy, we propose RocketKV, a two-stage KV cache compression method that combines permanent KV token eviction with dynamic KV token selection to accelerate the decode phase of LLM inference. In the first stage, we propose SnapKV++ for coarse-grain permanent KV token eviction on the input prompt. SnapKV++ improves upon SnapKV [20] with adaptive pooling size and full compatibility with grouped-query attention (GQA). In the second stage, RocketKV performs fine-grain dynamic KV token selection with a hybrid attention mechanism which estimates KV token indices with top-k attention scores by leveraging both head and sequence dimensional reductions. Combining these two stages together, RocketKV achieves significant memory bandwidth and storage savings with negligible accuracy loss across a wide variety of models and downstream tasks.
In summary, the contributions of the paper can be summarized as follows:
  • We analyze and identify the limitations of existing KV token dropping methods and then propose RocketKV with two-stage KV cache compression.
  • In the first stage, we propose SnapKV++ for prompt KV cache eviction that improves upon SnapKV by introducing adaptive pooling size and full compatibility with GQA. In the second stage, we propose hybrid attention for dynamic KV token selection at each decode step.
  • We conduct a comprehensive evaluation of RocketKV on a wide variety of models and downstream tasks. RocketKV consistently demonstrates comparable accuracy with full-KV attention while maintaining KV token budgets as low as 256.
  • We evaluate the efficiency of RocketKV on an NVIDIA H100 GPU and it shows up to 3$\times$ end-to-end speedup while reducing total memory consumption by up to 31% at the decode phase compared to the full KV cache.
2 Related Work
A number of recent approaches have focused on improving the efficiency of attention mechanisms in LLMs, particularly when dealing with long contexts. One feasible solution is to share KV cache across multiple layers [5] or selectively drop attention for some layers [21]. Others [29, 10] propose mixed attention designs where some layers use global attention while others use local attention. Multi-Query Attention (MQA) [26] and Grouped-Query Attention (GQA) [2] are widely adopted by many recent LLMs [1, 13, 32], which reduce KV cache by sharing them across multiple attention heads. All the above techniques directly modify the attention architecture and require integration since pre-training.
Another direction is to improve attention efficiency with training-free techniques. These techniques can be further categorized as prefill phase acceleration, decode phase acceleration, or accelerating both phases together. For example, StreamingLLM [13] combines initial and local-window attention to reduce the KV cache into a constant size regardless of the sequence length. DuoAttention [35] and RazorAttention [30] improve upon StreamingLLM by applying global attention on retrieval heads and StreamingLLM-style attention on rest heads. While these techniques can be applied to both the prefill and decode phases of LLM inference, other techniques focus on accelerating only one of the two phases.
For prefill phase attention acceleration, MInference [14] identifies three distinct patterns in long-context attention matrices that can be harnessed for efficient sparse operations with customized GPU kernels. SeerAttention [8] explores dynamic block-level sparsity in attention module with a learnable gate.
A common approach for attention acceleration at the decode phase is through permanent KV cache eviction, which can save both KV cache fetching bandwidth and storage requirements. H2O [37] observes that a small subset of tokens, known as heavy-hitters, dominates the attention computation, thus only keeps recent and heavy-hitter tokens. SnapKV [20] employs an observation window at the end of the prompt to identify critical KV tokens of the input prompt. Then it uses a clustering algorithm via pooling to retain critical KV token clusters without loss the completeness of the information. Quest [31] observes that permanent KV cache eviction could lead to inevitable accuracy loss, and proposes query-aware selection on top-k KV tokens based on approximation attention with representative vectors of contiguous key cache pages. On the other hand, SparQ and Loki [25, 28] conduct approximation attention by selecting only important indices on the head dimension instead. The above approaches can save KV cache compute and data fetching from memory but not KV cache storage. InfiniGen [18] tackles this challenge by offloading the entire KV cache to CPU memory and only fetching the selected KV tokens to GPU memory when needed. MagicPIG[6] discovers that using importance sampling is more efficient than top-k estimation and proposes an approximation attention solution leveraging locality sensitive hashing (LSH) and CPU offloading.
In contrast to prior work, RocketKV combines both permanent KV cache eviction and dynamic KV token selection as two consecutive stages for accelerating the decode phase and utilizes the advantages of both worlds. As a result, it achieves remarkable memory bandwidth and storage savings without the need for sophisticated system-level optimizations such as CPU offloading, while maintaining remarkable accuracy even at low token budgets.
3 Proposed Method: RocketKV
3.1 Observation
As discussed previously in Sections 1 and 2, existing training-free KV cache eviction methods for decode phase acceleration can be categorized into two types: permanent KV cache eviction and dynamic KV token selection. To understand the effectiveness of strategies in each type, Figure 1 presents the accuracy comparison of four different methods for qasper benchmark in LongBench [4] on Mistral-7B-Instruct-v0.2 model, where DuoAttention and SnapKV belong to the first type, and Quest and SparQ belong to the second type. As shown in the figure, the accuracy of all four methods drops significantly as the token budget becomes lower than 1024. In comparison, we present Exact-TopK as an oracle method that performs sparse attention on top-k KV token indices selected based on the exact attention scores. Exact-TopK achieves negligible accuracy drop compared to the Full-KV attention even with a token budget of 256, which indicates that all those practical methods fail to accurately predict top-k KV tokens at low token budgets.
To improve the prediction accuracy, we observe that permanent KV cache eviction and dynamic KV token selection are complementary to each other and can be seamlessly combined together. In fact, we can perform permanent KV cache eviction with a much larger token budget first and then conduct dynamic KV token selection on the remaining KV tokens. By doing this, the chance of permanently evicting important KV tokens is greatly reduced. Meanwhile, dynamic KV token selection only needs to perform on a subset of KV tokens rather than the original full set, significantly lowering the prediction difficulty.
Figure 1: Comparing permanent KV cache eviction and dynamic KV token selection for qasper benchmark.
3.2 RocketKV
Based on the above observation, we propose RocketKV, a two-stage KV cache compression method for decode phase acceleration. As shown in Figure 2, RocketKV performs coarse-grain KV cache eviction at the first stage. The purpose of this stage is to remove KV tokens with low importance while trying to keep the majority of important tokens. In the second stage, it conducts fine-grain dynamic KV token selection on the remaining KV tokens followed by top-k sparse attention. For a given token budget $t$ and associated compression ratio $c$, we evenly split the compression ratio across both stages so that each stage results in a compression ratio of $\sqrt{c}$. To achieve the best performance at each stage, we further propose a highly effective KV cache eviction and dynamic KV token selection method for each stage.
Figure 2: Overview of RocketKV with two consecutive stages.
3.3 First Stage: SnapKV++
Algorithm 1 SnapKV++
  Input: sequence length $S$, query vector $q$, key tensor $K$, value tensor $V$, kernel sizes $ks_{1}$ and $ks_{2}$, kernel threshold $Thr$, top-k KV indices $k$
  # compute Attention scores based on observation window
  $s_{1}\leftarrow score(q_{obs},K_{pre})$
  # accumulate attention score in seq. and group dim
  $s_{2}\leftarrow sum(s_{1})$
  # adaptively select kernel size for pooling
  $ks\leftarrow ks_{1}$ if $S\geq Thr$ else $ks_{2}$
  # perform pooling in seq. dim
  $s_{3}\leftarrow pool(s_{2},ks)$
  # get indices with top-k attention scores
  $i\leftarrow argtopk(s_{3},k)$
  # generate pruned KV cache
  $K_{cache}\leftarrow K_{[i,obs]},V_{cache}\leftarrow V_{[i,obs]}$
  return $K_{cache},V_{cache}$
Algorithm 2 Hybrid Attention (only contains step 2 and 3)
  Input:query vector $q$, key tensor $K$, value tensor $V$, element-wise max/min key tensor $K_{max}/K_{min}$, top-r $q$ indices $r$, top-k KV indices $k$
  # get top-r indices along head dim from sum of $|q|$ in group dim
  $i_{1}\leftarrow argtopk(sum(|q|,dim=group),r)$
  # get signs of top-r indices from sum of q in group dim
  $g\leftarrow sign(sum(q_{[i_{1}]},dim=group))$
  # fetch corresponding indices from paged min or max
  $P\leftarrow K_{max[i:g_{i}\geq 0]},K_{min[i:g_{i}<0]}$
  # compute approximation attention scores
  $s_{1}\leftarrow score(q_{[i1]},P)$
  # get indices with top-k attention scores along seq. dim
  $i_{2}\leftarrow argtopk(s_{1},k)$
  # perform sparse attention
  $y\leftarrow attn(q,K_{[i2]},V_{[i2]})$
  return $y$
Figure 3: Illustration of hybrid attention with figure (up) and algorithm (bottom).
In the first stage, we propose SnapKV++ as an efficient KV cache compression method on the input prompt. Compared with the original SnapKV, SnapKV++ introduces two enhancements that greatly improve the performance: 1) Full compatibility to GQA, and 2) adaptive kernel size for pooling.
The original SnapKV method keeps crucial KV tokens based on aggregated per-head attention scores across the observation window. In the case of GQA, each attention head within an attention group will keep a separate set of KV cache tokens, which could introduce redundant storage of the same KV token. To be fully compatible with GQA, SnapKV++ accumulates per-group rather than per-head attention score, and hence the selected KV tokens are shared across the entire attention group.
SnapKV uses pooling along the sequence dimension so that the critical KV tokens are selected along with their neighbor tokens. It demonstrates better accuracy with pooling because it retains the completeness of selected information. The employed kernel sizes for pooling are quite small (e.g., a kernel size of 7 for LongBench [4]). Since SnapKV++ is only used for coarse-grain KV token eviction at the first stage, we discovered that the optimal kernel sizes for pooling are much larger. In fact, the optimal kernel size can grow as the input sequence length increases. Therefore, SnapKV++ provides a simple adaptive kernel size selection mechanism based on the input sequence length: it selects a smaller kernel size if the input sequence length is shorter than a threshold value, otherwise, it will switch to a larger kernel size. Those two kernel sizes and threshold values are empirically determined in our work as shown in Section 4.4. We found this mechanism is sufficient to capture close-to-optimal performance in our studies, so we did not explore more sophisticated algorithms.
The overall algorithm of SnapKV++ is illustrated in Algorithm 1. We found that with those two enhancements, SnapKV++ greatly outperforms the original SnapKV method as shown in Section 4.4.
3.4 Second Stage: Hybrid Attention
Previous methods on dynamic KV token selection often leverage the sparsity in the head or sequence dimension of the Q or K tensor [31, 25, 28] to estimate top-k attention scores. Unfortunately, relying on the sparsity of a single dimension can only achieve a certain degree of compression ratio, beyond which the accuracy could drop rapidly, as shown previously in Figure 1. In contrast, we propose a hybrid attention method which takes advantage of sparsity in both head and sequence dimensions together to achieve better estimation accuracy on KV token indices with top-k attention scores.
Figure 3 shows the detailed implementation of our proposed algorithm, which is inspired by Quest [31] and SparQ [25]. Our hybrid attention algorithm can be decomposed into three steps:
  • Step 1: Group tokens in key tensor into consecutive pages along the sequence dimension and store element-wise min and max of each page as auxiliary storage. They are stored along the head dimension to enable efficient gathering in Step 2. The auxiliary storage is updated accordingly each time a new key token is generated and stored.
  • Step 2: For each query q, accumulate $|q|$ in group dimension and find the r largest indices along the head dimension. Then fetch only the corresponding indices in either element-wise min or max tensors depending on the sign of the sum of $q$ in group dimension at those indices. Finally calculate approximation attention scores based on those partial $q$ and element-wise vectors, and find top-k indices along the sequence dimension.
  • Step 3: Perform sparse attention by fetching the full key and value vectors from the top-k indices.
In our hybrid attention algorithm, the $K$ tensor is compressed along the sequence dimension first for auxiliary storage, and then selectively fetched along the head dimension for approximation attention calculation. For a given compression ratio $c$, we evenly split it between those two steps so that each step results in a compression ratio of $\sqrt{c}$. This would avoid significant accuracy loss caused by over-compression on a single dimension. Our method is fully compatible with GQA because all selections on the key and value vectors are based on per attention group rather than attention head, similar to our enhancement in SnapKV++.
Table 1: Normalized KV cache storage (including auxiliary data) and traffic comparison between RocketKV and other methods.
Method Compression Ratio Storage Traffic
SnapKV $c$ $1/c$ $1/c$
Quest $c$ $1+1/c$ $1/c$
SparQ $c$ 2 $1/c$
RocketKV $c$ $1/\sqrt{c}+2/c^{3/4}$ $1/c$
3.5 Cost Analysis and Comparison
The decode phase of LLM inference is typically memory bound [25] so the time spent in the attention module is roughly proportional to the total memory traffic. In this work, we use token budget $t$ to estimate the amount of memory traffic for each attention operation in the decode phase (we mainly focus on KV cache traffic since it contributes to the majority of memory traffic in this scenario). For example, a token budget of 512 means each attention module needs to fetch an equivalent total amount of 512 key and value pairs from the memory. For models with GQA, this token budget is defined for the entire attention group rather than each attention head. For a given sequence length $S$, the total compression ratio $c$ can be defined as $c=S/t$.
In RocketKV, the total compression ratio is evenly split between two stages with each stage having a compression ratio of $\sqrt{c}$. Therefore, in the first stage, SnapKV++ results in both KV cache storage and traffic reduction of $\sqrt{c}$. In the second stage, we need to take the additional memory storage overhead introduced by approximation attention into consideration. Since we further evenly split the compression ratio $\sqrt{c}$ into $c^{1/4}$ between step 1 and 2 in the hybrid attention algorithm, it introduces a memory storage overhead of $(1/\sqrt{c})\times(1/c^{1/4})\times 2=2/c^{3/4}$ where $1/\sqrt{c}$ is the relative KV cache storage after SnapKV++ and both element-wise min and max tensors will introduce a storage overhead of $1/c^{1/4}$ on top of it. Therefore, the total KV cache storage and traffic in RocketKV are $1/\sqrt{c}+2/c^{3/4}$ and $1/c$ of the full KV baseline, respectively. Table 1 compares the KV cache storage and traffic of RocketKV against other methods at a given compression ratio $c$ and we can see that while all methods lead to KV cache traffic savings, only RocketKV and SnapKV provide additional KV cache storage savings but Quest and SparQ require extra storage for auxiliary data.
In terms of system-level optimizations, RocketKV is fully compatible with FlashAttention [7] as it does not modify attention in the prefill phase, as well as Tensor Parallelism [27] as all operations are symmetric across attention heads/groups.
4 Experiments
4.1 Experimental Settings
(a) LongBench, Llama3.1-8B-Ins
(b) LongBench, Mistral-7B-Ins-v0.2
(c) LongBench, LongChat-7B-v1.5
(d) NIAH, Llama3.1-8B-Ins
(e) NIAH, Mistral-7B-Ins-v0.2
(f) NIAH, LongChat-7B-v1.5
(g) Llama3.1-8B-Ins, SeqLen=16K
(h) Llama3.1-8B-Ins, SeqLen=32K
(i) Llama3.1-8B-Ins, SeqLen=64K
(j) Llama3.1-8B-Ins, SeqLen=96K
(k) Mistral-7B-Ins-v0.2,SeqLen=8K
(l) Mistral-7B-Ins-v0.2,SeqLen=16K
(m) Mistral-7B-Ins-v0.2,SeqLen=24K
(n) Mistral-7B-Ins-v0.2,SeqLen=32K
(o) LongChat-7B-v1.5,SeqLen=8K
(p) LongChat-7B-v1.5,SeqLen=16K
(q) LongChat-7B-v1.5,SeqLen=24K
(r) LongChat-7B-v1.5,SeqLen=32K
Figure 4: Comparing the accuracy of RocketKV with other methods under LongBench ((a) to (c)), Needle-in-a-Haystack ((d) to (f)), and RULER with various sequence lengths ((g) to (r)).
We conduct our experiments on three widely used long-context models: Llama3.1-8B-Instruct [1], Mistral-7B-Instruct-v0.2 [13], and LongChat-7B-v1.5-32k [19]. We use the notations Llama3.1-8B-Ins, Mistral-7B-Ins-v0.2, LongChat-7B-v1.5 for them, respectively throughout the paper. The first two models are based on GQA, while the last one is based on MHA. For downstream tasks, we utilize LongBench [4], Needle-in-a-Haystack [16] and RULER [11] benchmarks. We compare RocketKV with several other methods: Full-KV, Exact-TopK, DuoAttention [35], SnapKV [20], Quest [31], and SparQ [25]. Exact-TopK serves as an oracle method for sparse attention with exact top-k KV token selection. We evaluate all methods across various KV token budgets per attention group except for Full-KV. RocketKV, Quest, and SparQ involve additional memory traffic for top-k approximation, which is converted into an equivalent KV token budget such that the total token budget precisely reflects the total amount of memory loads in the attention module. More details on our experimental settings can be found in Appendix A.
4.2 Accuracy Results
In our accuracy evaluation, we vary the token budget of each method from 256 to 4096 except for the Full-KV and then compare the average accuracy across all individual tasks for each benchmark. More detailed results with accuracy breakdown for individual tasks can be found in Appendix B.
LongBench Benchmark: The first row in Figure 4 shows the average score comparison of RocketKV with other methods on LongBench across all three models. Based on the figure, we can see that RocketKV consistently outperforms all other methods, especially with the lower token budgets. For Llama3.1-8B-Ins, RocketKV achieves almost no accuracy loss with a token budget of 512 and above, and only 1.1% average accuracy drop with a token budget of 256. RocketKV results in slightly higher accuracy losses for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5, which might be because Llama3.1-8B-Ins is better-trained than the other two models, making it more robust to sparse attention methods, as a similar trend can be found with Exact-TopK. SparQ performs well on Llama3.1-8B-Ins and LongChat-7B-v1.5 with a token budget of 1024 and above but underperforms all other methods on Mistral-7B-Ins-v0.2. SnapKV achieves relatively good accuracy with a token budget of 1024 and above across all three models, however, it cannot further reduce the token budget to 512 or below for tasks in LongBench with 512 maximum generated tokens. This is because it only compresses the KV cache of the input prompt but not the generated tokens, so the token budget cannot reach 512 or below for those tasks. Although we cannot plot the average accuracy score of SnapKV for those cases, the accuracy scores for the remaining tasks are available in Appendix B.3, and we can still see widening accuracy gaps between SnapKV and RocketKV for those tasks.
Needle-in-a-Haystack Benchmark: Presented in the second row of Figure 4, RocketKV achieves near the Full-KV accuracy across all models, even with a token budget of 512, and experiences only negligible accuracy loss with a token budget of 256. In contrast, all other methods suffer substantial accuracy drops. For example, the accuracy of SnapKV for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5 decreases by more than 20% compared to the Full-KV, even with a token budget of 1024. We present the heatmaps of RocketKV under various token budgets in Appendix B.2.
RULER Benchmark: For the RULER benchmark, we evaluate all methods across three models with varying sequence lengths as shown in Figure 4 (rows 3-5). Again, RocketKV shows robust performance and a clear advantage over all other methods across various token budget and sequence length settings. Overall, we can see that the accuracy loss of RocketKV is negligible under short sequence lengths, and gradually becomes larger as sequence length increases. We believe this is because those models are less robust to sparse attention under longer sequence lengths. Notice that the accuracy gaps between RocketKV and other methods become even wider under longer sequence lengths.
(a) End-to-end speedup
(b) Peak memory saving
Figure 5: End-to-end speedup and peak memory savings of RocketKV with various token budgets compared to the Full-KV.
4.3 Efficiency Results
Our efficiency experiments are conducted with Llama3.1-8B-Ins [1] under FP16 precision, running on an NVIDIA H100 GPU at a batch size of 1. Similar to SparQ, we leverage gpt-fast [17], a low-latency Python-native
LLM framework for running the efficiency experiments. We found a python-based implementation of RocketKV under gpt-fast is sufficient to demonstrate its efficiency benefit, but it could be further improved with customized CUDA kernels and more advanced frameworks such as FlashInfer [36]. Figure 5 demonstrates the end-to-end speedups and peak memory savings of RocketKV with varying token budgets at the decode phase, with all values normalized to the Full-KV. Notice that the peak memory usage measured in the figure includes memory allocations for weights, activations, KV cache, and all other metadata at the decode phase. As shown in the figure, RocketKV with token budgets of 512 and 256 achieves up to 2.9$\times$ and 3$\times$ end-to-end speedups as well as 30% and 31.4% peak memory savings, respectively. Additionally, increasing the sequence lengths results in almost linear improvements in both speedup and peak memory savings.
(a) LongBench, Llama3.1-8B-Ins
(b) LongBench, Mistral-7B-Ins-v0.2
(c) NIAH, Llama3.1-8B-Ins
(d) NIAH, Mistral-7B-Ins-v0.2
Figure 6: LongBench and Needle-in-a-Haystack accuracy comparison between SnapKV and SnapKV with GQA enhancement.
4.4 Ablation Study
In this subsection, we present a series of ablation studies to further demonstrate the effectiveness of RocketKV. More ablation results can be found in Appendix B.
4.4.1 SnapKV with GQA enhancement
As illustrated in Figures 6, SnapKV with the grouped-query attention (GQA) [2] enhancement consistently outperforms the original SnapKV, especially with low token budgets. This is because with the GQA enhancement, the compressed KV tokens are shared across the entire attention group which is far more effective than storing them separately within each attention head. Recall that, the GQA enhancement does not make any difference for LongChat-7B-v1.5 since it uses MHA in the attention module.
(a) Llama3.1-8B-Ins, SeqLen=16K
(b) Llama3.1-8B-Ins,SeqLen=32K
(c) Llama3.1-8B-Ins, SeqLen=64K
(d) Llama3.1-8B-Ins, SeqLen=96K
(e) Mistral-7B-Ins-v0.2, SeqLen=8K
(f) Mistral-7B-Ins-v0.2,SeqLen=16K
(g) Mistral-7B-Ins-v0.2,SeqLen=24K
(h) Mistral-7B-Ins-v0.2,SeqLen=32K
(i) LongChat-7B-v1.5, SeqLen=8K
(j) LongChat-7B-v1.5, SeqLen=16K
(k) LongChat-7B-v1.5, SeqLen=24K
(l) LongChat-7B-v1.5, SeqLen=32K
Figure 7: Comparing RocketKV with different kernel sizes (7, 63, and 511) under RULER with various sequence lengths.
4.4.2 RocketKV kernel size selection
We run extensive experiments to examine the impact of the kernel sizes for pooling on the overall accuracy of RocketKV. We primarily focus on the RULER benchmark to understand the optimal kernel size selection across different sequence lengths, as shown in Figures 7. For Llama3.1-8B-Ins, a kernel size of 63 performs best for sequence lengths of 16K and 32K, while a kernel size of 511 performs best for sequence lengths of 64K and 96K. Therefore, we simply set the switching threshold to 48K as the middle point between 32K and 64K. For Mistral-7B-Ins-v0.2, a kernel size of 63 generally performs best across all sequence lengths so there is no need to switch to other kernel sizes. For LongChat-7B-v1.5, a kernel size of 63 performs best for a sequence length of 8K, and a kernel size of 511 performs best for remaining sequence lengths; so, we set the switching threshold to 12K. We found adaptive kernel selection with empirically determined kernel sizes and threshold generalizes well in other benchmarks (shown in Appendix B.1). Hence, we do not explore more sophisticated solutions.
(a) LongBench, Llama3.1-8B-Ins
(b) LongBench, Mistral-7B-Ins-v0.2
(c) LongBench, LongChat-7B-v1.5
(d) NIAH, Llama3.1-8B-Ins
(e) NIAH, Mistral-7B-Ins-v0.2
(f) NIAH, LongChat-7B-v1.5
Figure 8: LongBench and Needle-in-a-Haystack accuracy comparison among Hybrid, Quest, and SparQ.
4.4.3 Comparing Hybrid, Quest, and SparQ
To illustrate the effectiveness of hybrid attention, we compare the accuracy of the standalone hybrid attention mechanism (Hybrid) against Quest and SparQ in the Figures 8. In all cases, hybrid attention consistently outperforms Quest and SparQ, especially at low token budgets. This clearly demonstrates the advantage of hybrid attention which intelligently leverages approximations in both head and sequence dimensions compared to single dimension approximation methods such as Quest and SparQ.
5 Conclusion
RocketKV presents a novel, training-free approach to KV cache compression, addressing the challenges of memory bandwidth and capacity demands during the decode phase of LLM inference. RocketKV contains two consecutive stages: SnapKV++ for coarse-grain KV cache eviction and hybrid attention for fine-grain dynamic KV token selection. Our evaluations on models and long-context benchmarks demonstrate that RocketKV maintains comparable accuracy to full KV cache attention while significantly lowering memory bandwidth and capacity usage. At low token budgets of 256 or 512, RocketKV outperforms all other techniques by a considerable margin. It achieves up to 3$\times$ speedup and 31% peak memory reduction at the decode phase on an NVIDIA H100 GPU compared to the full KV cache baseline, highlighting its efficiency and potential for widespread application in optimizing LLM performance.
Appendix A Detailed Experiment Settings
A.1 Models
We evaluate our methods on the three widely used long-context models: Llama3.1-8B-Ins [1], Mistral-7B-Ins-v0.2 [13], and LongChat-7B-v1.5 [4]. They are all decoder-only transformer models [33]. Llama3.1-8B-Ins and Mistral-7B-Ins-v0.2 use GQA, while LongChat-7B-v1.5 uses MHA in the attention module. In terms of sequence length, Llama3.1-8B-Ins supports a maximum sequence length of 128K while the other two support up to 32K sequence length. Prior work [31, 6] usually skips the first two layers during KV cache compression to maintain accuracy. We found this is only necessary for LongChat-7B-v1.5; so, we conduct KV cache compression on all attention layers for Llama3.1-8B-Ins and Mistral-7B-Ins-v0.2 but skip the first two layers for LongChat-7B-v1.5.
A.2 Benchmarks
Our LongBench and Needle-in-a-Haystack settings are mostly following the prior work [34]. For LongBench, we set the maximum prompt length to 127,500 for Llama3.1-8B-Ins and 31,500 for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5. Prompts beyond the maximum prompt length will be middle-truncated by keeping the first and last half of the input tokens. For Needle-in-a-Haystack, we evaluate with 10 different input sequence lengths uniformly spanning from 2048 to 81,920 words and 10 different depths for each sequence length in Llama3.1-8B-Ins. The maximum word count of 81,920 roughly converts to 109K tokens by the tokenizer in Llama3.1-8B-Ins. For the other two models, the input sequence lengths range from 512 to 20480 words since they support up to 32K sequence length, which convert up to  30K and 31K tokens for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5, respectively. For RULER, we mostly follow the configurations in the original benchmark except for reducing the number of examples per task from 500 to 50 to speed up the evaluation.
A.3 Baselines
Since RocketKV primarily focuses on accelerating the decode phase of LLM inference, we do not perform any KV cache compression or sparse attention mechanism in the prefill phase across all comparing baselines including DuoAttention for fair comparison.
Exact-TopK: Exact-TopK serves as an oracle method to demonstrate the effectiveness of top-k sparse attention. We assume the top-k KV tokens can be directly identified with no cost so a given token budget will directly correspond to its top-k value.
DuoAttention: We set the initial and recent tokens to 128 and 512 respectively for streaming heads if the token budget is larger than 640, and vary the ratio of retrieval heads to match with the given token budget on average. If the token budget is lower than 640, all attention heads become streaming heads. In this case, we set the number of initial tokens for attention sink to 128 if the token budget is larger than 256, otherwise it is set to 20% of the token budget. The rest of the token budget is assigned to recent tokens within the sliding window.
SnapKV: We set the observation window size to 32 and kernel size for pooling to 7 following the SnapKV paper. Since SnapKV only compresses the KV cache of the input prompt, the token budget for the input prompt is calculated by deducting the number of maximum generated tokens from the total token budget. This budget is further divided by the number of heads per attention group in case of GQA as SnapKV performs on each attention head separately and the pruned KV cache is not shared within the attention group.
Quest: We evenly split the token budget between approximate attention for identifying top-k KV token indices and top-k sparse attention so that the token budget can accurately reflect total memory fetching bandwidth required in the attention module. The original Quest method is not compatible with GQA. We modify it to select indices of top-k accumulated attention scores per attention group rather than per attention head similar to our enhancement in SnapKV++.
SparQ: Similar to Quest, we evenly split the token budget between approximate attention for identifying top-k KV token indices and top-k sparse attention.
RocketKV: We evenly split the compression ratio between SnapKV++ in the first stage and hybrid attention in the second stage. For SnapKV++, we use an observation window size of 32 and adaptively select the kernel size for pooling between 63 and 511, depending on the input sequence length. For Llama3.1-8B-Ins, we use the kernel size of 63 if the input sequence length is shorter than 48K, otherwise 511 is used. For LongChat-7B-v1.5, the input sequence length threshold is 12K, and Mistral-7B-Ins-v0.2 always selects a kernel size of 63. The adaptive selection policy is empirically determined based on our ablation study in Section 4.4. For hybrid attention, we evenly split the token budget between approximate attention for identifying top-k indices and top-k sparse attention.
Appendix B Detailed Results
B.1 Further Ablation Study
This subsection presents additional results from the ablation studies discussed in the main paper. These results further enhance our observations in Section 4.4.
(a) Llama3.1-8B-Ins, SeqLen= 16K
(b) Llama3.1-8B-Ins, SeqLen=32K
(c) Llama3.1-8B-Ins, SeqLen=64K
(d) Llama3.1-8B-Ins, SeqLen=96K
(e) Mistral-7B-Ins-v0.2, SeqLen=8K
(f) Mistral-7B-Ins-v0.2, SeqLen=16K
(g) Mistral-7B-Ins-v0.2, SeqLen=24K
(h) Mistral-7B-Ins-v0.2, SeqLen=32K
Figure 9: RULER accuracy comparison between SnapKV and SnapKV with GQA enhancement.
(a) LongBench, Llama3.1-8B-Ins
(b) LongBench, Mistral-7B-Ins-v0.2
(c) LongBench, LongChat-7B-v1.5
(d) NIAH, Llama3.1-8B-Ins
(e) NIAH, Mistral-7B-Ins-v0.2
(f) NIAH, LongChat-7B-v1.5
Figure 10: Comparing RocketKV with different kernel sizes (7, 63 and 511) under LongBench and Needle-in-a-Haystack
(a) Llama3.1-8B-Ins, SeqLen = 16K
(b) Llama3.1-8B-Ins, SeqLen= 32K
(c) Llama3.1-8B-Ins, SeqLen=64K
(d) Llama3.1-8B-Ins, SeqLen=96K
(e) Mistral-7B-Ins-v0.2, SeqLen=8K
(f) Mistral-7B-Ins-v0.2, SeqLen=16K
(g) Mistral-7B-Ins-v0.2, SeqLen=24K
(h) Mistral-7B-Ins-v0.2, SeqLen=32K
(i) LongChat-7B-v1.5, SeqLen=8K
(j) LongChat-7B-v1.5, SeqLen=16K
(k) LongChat-7B-v1.5, SeqLen=24K
(l) LongChat-7B-v1.5, SeqLen=32K
Figure 11: accuracy comparison among Hybrid, Quest, and SparQ under RULER with various sequence lengths
B.2 Needle-in-a-Haystack Visualization
Needle-in-a-haystack (NIAH) is a type of synthetic challenge designed to test how effectively an LLM can retrieve specific information in a large volume of text [16]. In Figures 1213,  14, the x-axis shows the document length (i.e., ”haystack”), while the y-axis marks the relative position of the ”needle” (i.e., a short sentence) within the text. As the results show, RocketKV can retrieve the needle with almost the same accuracy as the Full-KV throughout the whole text across all three models with token budgets as low as 512, and only incurs small accuracy losses with a token budget of 256.
(a) Token Budget = 256
(b) Token Budget = 512
(c) Token Budget = 1024
(d) Token Budget = 2048
(e) Token Budget = 4096
(f) Full-KV
Figure 12: Needle-in-a-Haystack visualization results of RocketKV on Llama3.1-8B-Ins.
(a) Token Budget = 256
(b) Token Budget = 512
(c) Token Budget = 1024
(d) Token Budget = 2048
(e) Token Budget = 4096
(f) Full-KV
Figure 13: Needle-in-a-Haystack visualization results of RocketKV on Mistral-7B-Ins-v0.2.
(a) Token Budget = 256
(b) Token Budget = 512
(c) Token Budget = 1024
(d) Token Budget = 2048
(e) Token Budget = 4096
(f) Full-KV
Figure 14: Needle-in-a-Haystack visualization results of RocketKV on LongChat-7B-v1.5.
B.3 Detailed Accuracy Results
In this section, we present 18 detailed tables that display the accuracy results discussed in previous sections. These tables offer additional information and insights into the results we captured.
Table 2: Average results of LongBench and NIAH for Llama3.1-8B-Ins
Token Budget
Method
Single QA
Multi. QA
Summ.
Few-shot
Synthetic
Code
LB Avg.
NIAH(%)
N/A Full-KV 43.5 44.5 29.2 69.3 53.9 62.0 52.2 100.0
256 Exact-TopK 43.6 44.4 28.9 69.5 53.6 65.3 52.6 100.0
DuoAttention 24.4 33.5 18.2 39.0 51.2 60.1 37.6 20.0
SnapKV 12.3 35.8 N/A 30.8 53.5 54.4 N/A 0.0
Quest 16.0 14.0 14.3 14.4 4.0 41.4 17.8 10.7
SparQ 26.3 18.7 18.2 24.5 15.5 47.0 25.7 10.3
Hybrid 41.3 41.9 26.8 64.2 51.7 61.1 49.5 44.3
RocketKV 43.4 43.0 24.9 67.3 55.6 65.6 51.1 99.0
512 Exact-TopK 43.7 44.5 28.8 69.1 53.7 63.1 52.2 100.0
DuoAttention 27.7 34.9 21.6 47.4 51.5 61.7 41.2 21.0
SnapKV 33.9 41.0 N/A 50.4 53.3 59.7 N/A 99.0
Quest 27.9 27.7 19.8 38.0 43.5 52.6 35.1 18.0
SparQ 37.1 35.1 24.5 53.2 48.0 57.3 43.8 14.7
Hybrid 43.8 43.2 28.5 68.6 53.1 65.2 52.1 73.7
RocketKV 43.9 44.7 26.3 69.6 53.7 65.3 52.2 99.7
1024 Exact-TopK 43.8 44.6 29.0 69.3 53.5 61.9 52.2 100.0
DuoAttention 31.8 36.4 23.7 54.8 53.3 64.2 44.5 26.3
SnapKV 36.2 41.8 21.5 57.9 54.8 60.6 46.2 99.7
Quest 36.5 37.5 24.8 54.9 50.9 57.3 45.0 37.3
SparQ 41.5 42.9 28.6 67.8 52.0 62.2 51.1 26.0
Hybrid 43.9 44.0 28.8 69.7 54.0 64.6 52.5 94.3
RocketKV 43.5 44.0 27.8 68.7 54.0 65.2 52.1 100.0
2048 Exact-TopK 43.6 44.5 29.2 69.5 53.8 61.1 52.1 100.0
DuoAttention 39.3 40.9 25.4 64.5 52.2 64.8 49.2 28.3
SnapKV 40.1 42.7 24.1 63.3 53.3 63.4 49.1 99.7
Quest 42.5 43.0 28.4 65.5 52.3 59.8 50.5 64.3
SparQ 43.5 43.5 29.3 69.4 53.6 62.9 52.1 56.3
Hybrid 43.7 43.8 29.2 69.4 53.6 62.3 52.2 100.0
RocketKV 43.9 43.9 28.6 68.7 53.7 64.2 52.2 100.0
4096 Exact-TopK 43.5 44.8 29.1 69.3 53.3 60.4 52.0 100.0
DuoAttention 42.4 41.7 27.5 68.9 53.2 63.9 51.2 36.7
SnapKV 42.5 44.0 26.1 67.4 53.1 63.2 51.0 100.0
Quest 43.9 44.3 28.9 69.0 53.5 60.0 51.9 94.3
SparQ 43.7 44.0 29.1 69.2 53.1 63.2 52.3 89.3
Hybrid 43.9 44.1 29.1 69.4 53.5 61.1 52.1 100.0
RocketKV 43.5 43.7 28.9 69.7 53.5 62.9 52.2 100.0
Table 3: Individual results of LongBench for Llama3.1-8B-Ins
Token Budget
Method
narrativeqa
qasper
mul.fieldqa
hotpotqa
2wikimqa
musique
gov-report
qmsum
multi-news
trec
triviaqa
samsum
pass-ret.
lcc
repobench-p
passagecnt
N/A Full-KV 30.2 45.5 54.9 55.5 46.7 31.3 35.2 25.2 27.2 72.5 91.7 43.8 99.5 65.1 58.8 8.4
256 Exact-TopK 30.7 44.7 55.4 55.0 46.5 31.7 34.8 25.1 26.8 71.5 92.2 44.8 99.5 67.1 63.6 7.8
DuoAttention 19.4 22.8 31.1 42.3 35.9 22.4 16.9 20.6 17.2 20.0 57.4 39.7 98.0 61.2 59.1 4.5
SnapKV 0.0 0.0 36.8 47.5 37.8 22.2 N/A N/A N/A 17.5 75.0 0.0 99.0 55.8 53.0 8.0
Quest 3.0 22.4 22.6 14.9 23.2 4.1 9.8 12.7 20.5 11.5 17.3 14.4 7.5 49.5 33.3 0.5
SparQ 5.0 38.0 35.8 19.7 28.8 7.5 13.7 15.0 25.8 16.5 32.1 24.8 28.8 60.1 34.0 2.2
Hybrid 26.4 43.5 54.1 54.2 43.7 27.8 30.7 23.2 26.6 63.5 87.0 42.2 97.0 66.3 56.0 6.5
RocketKV 30.3 43.5 56.4 54.9 45.4 28.8 26.7 23.8 24.1 67.5 91.0 43.5 99.5 66.7 64.4 11.6
512 Exact-TopK 30.9 45.0 55.1 55.3 47.0 31.0 34.5 25.0 26.9 71.5 92.0 43.9 99.5 65.0 61.1 7.9
DuoAttention 23.0 25.0 35.1 44.8 36.4 23.6 22.3 20.2 22.4 34.5 65.9 41.7 99.5 63.0 60.4 3.5
SnapKV 25.6 28.2 47.8 52.9 43.0 27.2 N/A N/A N/A 27.0 85.3 38.8 97.5 60.7 58.7 9.2
Quest 9.0 36.5 38.2 32.2 36.4 14.4 16.1 17.7 25.6 35.2 51.2 27.5 82.0 59.7 45.4 5.0
SparQ 13.6 45.4 52.1 43.8 39.4 22.2 25.6 21.2 26.7 49.5 72.2 37.9 92.0 64.2 50.5 4.1
Hybrid 30.4 45.2 55.8 53.8 46.1 29.8 33.8 24.9 26.9 70.5 90.1 45.2 99.0 66.3 64.2 7.1
RocketKV 30.4 45.0 56.1 55.4 47.6 31.2 29.5 24.6 24.8 73.0 91.7 44.0 99.5 66.2 64.5 7.8
1024 Exact-TopK 30.8 45.5 55.0 54.7 48.0 31.1 34.7 25.1 27.1 72.5 91.5 43.8 99.5 64.2 59.6 7.5
DuoAttention 24.1 30.4 41.1 45.3 39.4 24.4 24.5 21.2 25.3 45.5 76.3 42.5 98.5 66.8 61.5 8.0
SnapKV 26.5 33.1 49.1 54.4 43.3 27.7 20.7 23.1 20.6 43.0 90.0 40.5 99.5 62.6 58.6 10.0
Quest 15.4 44.4 49.8 48.9 40.8 22.9 27.1 21.0 26.4 57.0 70.0 37.9 99.0 62.6 52.0 2.7
SparQ 23.1 46.1 55.1 54.7 42.9 31.2 34.0 24.9 27.0 69.5 89.9 44.0 99.5 65.0 59.3 4.5
Hybrid 31.1 45.3 55.3 54.2 46.5 31.2 34.6 24.9 26.9 72.5 92.1 44.5 99.5 66.3 62.9 8.5
RocketKV 30.1 44.4 56.2 54.6 47.0 30.5 32.0 25.2 26.2 71.0 91.8 43.2 99.5 66.6 63.8 8.4
2048 Exact-TopK 30.7 45.3 54.7 54.8 47.2 31.3 34.9 25.4 27.2 73.0 91.5 44.1 99.5 63.8 58.4 8.2
DuoAttention 25.6 42.3 50.0 52.3 42.1 28.1 27.5 22.1 26.6 66.0 84.9 42.8 98.5 66.1 63.6 6.0
SnapKV 28.5 39.1 52.7 53.4 46.5 28.4 24.6 23.9 23.7 55.5 91.7 42.6 99.5 65.5 61.3 7.2
Quest 28.0 45.5 54.2 54.3 44.7 29.8 33.1 24.9 27.2 68.5 84.5 43.4 99.5 63.2 56.5 5.1
SparQ 30.3 45.2 54.9 53.7 46.5 30.4 35.3 25.4 27.1 73.0 91.2 44.0 99.5 65.3 60.5 7.7
Hybrid 30.7 45.2 55.2 54.1 46.2 31.1 35.1 25.1 27.4 73.0 91.2 44.2 99.5 64.3 60.3 7.7
RocketKV 30.8 45.4 55.5 54.7 47.0 30.0 33.3 25.5 27.1 71.0 91.5 43.7 99.5 66.3 62.1 7.9
4096 Exact-TopK 30.2 45.5 54.8 55.5 47.1 31.8 35.2 25.1 27.2 72.5 91.7 43.7 99.5 63.4 57.5 7.2
DuoAttention 26.5 45.1 55.5 55.0 43.3 26.7 31.9 23.7 27.0 73.5 89.9 43.3 99.5 65.5 62.3 7.0
SnapKV 29.4 44.2 53.9 55.8 46.0 30.1 28.2 24.3 25.7 67.0 92.0 43.1 99.5 65.1 61.2 6.8
Quest 30.7 45.5 55.4 55.6 46.0 31.3 34.9 24.8 27.1 72.0 90.9 44.0 99.5 63.3 56.8 7.4
SparQ 30.8 45.3 54.9 54.7 46.4 31.0 34.9 25.2 27.2 72.5 91.5 43.7 99.5 65.6 60.7 6.6
Hybrid 31.2 45.7 54.7 54.3 46.4 31.6 34.9 25.3 27.2 72.5 91.5 44.2 99.5 63.7 58.6 7.5
RocketKV 30.4 45.3 54.9 54.2 46.4 30.5 34.2 25.3 27.3 74.0 91.5 43.7 99.5 65.7 60.2 7.5
Table 4: Results of RULER, SeqLen=16000 for Llama3.1-8B-Ins
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.2 60.6 92.7 84.0 50.0 91.3
256 Exact-TopK 100.0 100.0 100.0 100.0 100.0 98.0 99.0 100.0 94.8 65.8 87.3 80.0 50.0 90.4
DuoAttention 0.0 2.0 2.0 2.0 0.0 0.0 6.0 5.0 1.6 2.8 67.3 42.0 28.0 12.2
SnapKV 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 62.0 42.0 8.0
Quest 2.0 4.0 0.0 12.0 0.0 0.0 1.0 2.5 0.0 0.6 33.3 6.0 8.0 5.3
SparQ 4.0 0.0 0.0 0.0 2.0 0.0 1.0 0.0 5.2 5.6 53.3 26.0 16.0 8.7
Hybrid 100.0 96.0 74.0 96.0 66.0 0.0 80.5 85.0 82.8 11.8 80.7 80.0 44.0 69.0
RocketKV 100.0 100.0 100.0 100.0 100.0 98.0 98.0 100.0 88.8 13.4 90.7 82.0 52.0 86.4
512 Exact-TopK 100.0 100.0 100.0 100.0 100.0 100.0 99.0 100.0 96.8 69.8 86.7 80.0 50.0 90.9
DuoAttention 2.0 2.0 4.0 2.0 4.0 0.0 6.0 5.0 3.2 0.4 90.0 42.0 20.0 13.9
SnapKV 100.0 60.0 0.0 34.0 26.0 0.0 2.0 2.5 3.6 0.0 53.3 68.0 40.0 30.0
Quest 98.0 74.0 0.0 100.0 26.0 0.0 22.0 16.0 60.4 4.0 52.7 34.0 22.0 39.2
SparQ 60.0 54.0 0.0 28.0 2.0 0.0 23.0 11.0 21.2 13.0 73.3 36.0 30.0 27.0
Hybrid 100.0 100.0 98.0 100.0 90.0 66.0 95.0 99.5 90.4 24.0 88.7 82.0 52.0 83.5
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 99.5 100.0 93.6 27.4 93.3 82.0 50.0 88.1
1024 Exact-TopK 100.0 100.0 100.0 100.0 100.0 100.0 99.5 100.0 98.8 71.4 89.3 80.0 50.0 91.5
DuoAttention 2.0 2.0 4.0 2.0 6.0 2.0 6.0 4.5 2.8 0.0 90.0 34.0 22.0 13.6
SnapKV 100.0 92.0 0.0 76.0 60.0 0.0 36.5 55.5 69.6 1.6 82.7 80.0 46.0 53.8
Quest 94.0 100.0 38.0 100.0 84.0 0.0 66.5 79.5 80.4 9.2 70.0 68.0 50.0 64.6
SparQ 100.0 100.0 100.0 100.0 98.0 54.0 97.0 99.0 85.2 66.0 86.7 84.0 54.0 86.5
Hybrid 100.0 100.0 100.0 100.0 100.0 98.0 99.5 100.0 94.4 51.8 88.0 82.0 50.0 89.5
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 94.8 47.2 96.7 80.0 54.0 90.2
2048 Exact-TopK 100.0 100.0 100.0 100.0 100.0 100.0 99.5 100.0 99.2 64.2 90.0 82.0 50.0 91.1
DuoAttention 4.0 20.0 6.0 20.0 24.0 4.0 6.0 8.0 2.4 2.6 89.3 52.0 34.0 20.9
SnapKV 100.0 96.0 0.0 92.0 84.0 2.0 59.0 82.0 92.0 11.6 87.3 76.0 48.0 63.8
Quest 94.0 96.0 60.0 98.0 92.0 28.0 91.5 90.5 80.8 19.8 86.7 80.0 50.0 74.4
SparQ 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.2 77.2 88.0 82.0 50.0 92.0
Hybrid 100.0 100.0 100.0 100.0 100.0 100.0 99.5 100.0 97.6 66.8 86.7 80.0 50.0 90.8
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.0 59.0 96.7 80.0 50.0 90.9
4096 Exact-TopK 100.0 100.0 100.0 100.0 100.0 100.0 98.5 100.0 99.2 57.4 90.7 84.0 50.0 90.8
DuoAttention 20.0 88.0 86.0 98.0 100.0 46.0 71.5 77.0 58.0 14.2 88.0 82.0 52.0 67.7
SnapKV 100.0 98.0 12.0 96.0 96.0 6.0 75.5 94.0 94.8 22.8 92.0 78.0 54.0 70.7
Quest 100.0 94.0 96.0 100.0 96.0 98.0 98.0 98.0 98.8 54.6 88.0 82.0 50.0 88.7
SparQ 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.2 75.2 92.0 82.0 50.0 92.2
Hybrid 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.2 77.0 91.3 82.0 50.0 92.3
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.2 59.6 94.7 82.0 50.0 91.2
Table 5: Results of RULER, SeqLen=32000 for Llama3.1-8B-Ins
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 100.0 100.0 100.0 100.0 100.0 99.0 99.5 96.4 6.0 94.0 74.0 54.0 86.4
256 Exact-TopK 100.0 100.0 100.0 100.0 100.0 90.0 99.0 99.5 91.6 24.6 81.3 76.0 54.0 85.8
DuoAttention 0.0 4.0 4.0 4.0 2.0 0.0 4.5 4.5 2.4 1.0 52.0 38.0 32.0 11.4
SnapKV 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 46.0 44.0 6.9
Quest 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.2 22.0 0.0 0.0 1.8
SparQ 44.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 2.8 0.4 18.7 4.0 8.0 6.0
Hybrid 98.0 96.0 38.0 94.0 16.0 0.0 52.0 57.5 70.4 0.4 48.0 62.0 38.0 51.6
RocketKV 100.0 100.0 100.0 100.0 100.0 94.0 95.5 97.5 83.6 1.0 73.3 82.0 54.0 83.1
512 Exact-TopK 100.0 100.0 100.0 100.0 100.0 100.0 99.0 99.0 90.8 17.2 82.0 78.0 52.0 86.0
DuoAttention 0.0 4.0 8.0 4.0 2.0 0.0 4.5 4.5 3.6 0.0 70.7 40.0 30.0 13.2
SnapKV 100.0 72.0 0.0 22.0 28.0 0.0 3.0 1.0 3.2 0.0 21.3 68.0 44.0 27.9
Quest 100.0 26.0 16.0 4.0 2.0 0.0 1.5 1.5 7.2 1.2 45.3 10.0 16.0 17.7
SparQ 62.0 0.0 0.0 0.0 0.0 0.0 0.5 0.5 3.2 1.6 26.7 24.0 26.0 11.1
Hybrid 100.0 96.0 90.0 100.0 72.0 6.0 89.0 96.0 83.2 1.2 60.7 80.0 52.0 71.2
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 97.0 99.5 89.2 5.4 81.3 80.0 52.0 85.0
1024 Exact-TopK 100.0 100.0 100.0 100.0 100.0 100.0 99.0 98.5 91.6 10.6 87.3 74.0 52.0 85.6
DuoAttention 0.0 4.0 8.0 4.0 2.0 0.0 4.5 4.0 4.0 0.0 71.3 42.0 34.0 13.7
SnapKV 100.0 88.0 0.0 74.0 58.0 0.0 33.5 61.0 65.2 0.4 56.0 68.0 50.0 50.3
Quest 90.0 78.0 18.0 76.0 46.0 0.0 34.0 31.0 38.8 1.0 50.0 20.0 22.0 38.8
SparQ 84.0 54.0 0.0 46.0 8.0 0.0 27.0 23.0 21.6 4.2 53.3 44.0 38.0 31.0
Hybrid 100.0 100.0 100.0 100.0 100.0 82.0 97.5 99.5 88.0 6.0 72.7 80.0 58.0 83.4
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 98.0 100.0 91.6 8.0 80.0 78.0 54.0 85.4
2048 Exact-TopK 100.0 100.0 100.0 100.0 100.0 100.0 99.0 98.5 95.6 7.0 90.7 74.0 54.0 86.1
DuoAttention 0.0 4.0 8.0 6.0 8.0 0.0 4.5 3.5 3.2 0.0 72.0 40.0 36.0 14.2
SnapKV 100.0 96.0 0.0 90.0 70.0 0.0 67.0 84.0 85.6 5.8 58.7 66.0 48.0 59.3
Quest 86.0 100.0 6.0 86.0 76.0 0.0 43.5 66.0 54.4 2.2 68.0 60.0 42.0 53.1
SparQ 100.0 100.0 98.0 98.0 96.0 36.0 98.0 98.0 82.0 34.2 74.7 76.0 58.0 80.7
Hybrid 100.0 100.0 100.0 100.0 100.0 100.0 99.0 99.0 92.0 14.6 76.7 76.0 56.0 85.6
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 99.5 100.0 93.2 18.2 82.0 76.0 52.0 86.2
4096 Exact-TopK 100.0 100.0 100.0 100.0 100.0 100.0 99.0 98.5 96.8 7.0 93.3 74.0 54.0 86.4
DuoAttention 4.0 26.0 14.0 36.0 16.0 2.0 5.5 10.5 3.2 0.0 76.7 50.0 48.0 22.5
SnapKV 100.0 96.0 14.0 98.0 80.0 2.0 76.0 93.5 88.8 9.2 68.7 74.0 54.0 65.7
Quest 82.0 100.0 70.0 90.0 92.0 38.0 87.0 87.5 88.4 5.6 78.7 78.0 52.0 73.0
SparQ 100.0 100.0 100.0 100.0 100.0 100.0 99.5 99.0 95.2 20.6 86.7 74.0 52.0 86.7
Hybrid 100.0 100.0 100.0 100.0 100.0 100.0 99.5 99.0 93.2 19.8 79.3 74.0 52.0 85.9
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.0 93.6 15.8 90.0 76.0 52.0 86.6
Table 6: Results of RULER, SeqLen=64000 for Llama3.1-8B-Ins
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 100.0 100.0 100.0 100.0 100.0 95.5 97.5 94.8 0.0 83.3 76.0 56.0 84.9
256 Exact-TopK 100.0 100.0 100.0 100.0 100.0 58.0 94.5 97.5 93.6 0.4 63.3 72.0 56.0 79.6
DuoAttention 0.0 2.0 0.0 8.0 0.0 0.0 5.0 3.0 0.0 0.4 84.0 44.0 32.0 13.7
SnapKV 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 52.0 32.0 6.5
Quest 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.4 25.3 0.0 2.0 2.1
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Hybrid 100.0 24.0 0.0 28.0 2.0 0.0 4.5 6.0 48.8 0.2 22.7 36.0 24.0 22.8
RocketKV 100.0 100.0 100.0 100.0 92.0 84.0 83.5 94.0 79.6 0.0 78.7 70.0 48.0 79.2
512 Exact-TopK 100.0 100.0 100.0 100.0 100.0 80.0 98.0 98.0 92.8 0.4 67.3 76.0 54.0 82.0
DuoAttention 0.0 2.0 0.0 8.0 0.0 0.0 5.0 3.5 0.4 0.8 96.0 38.0 28.0 14.0
SnapKV 100.0 60.0 0.0 28.0 12.0 0.0 1.0 0.0 3.2 0.0 20.0 62.0 48.0 25.7
Quest 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.5 0.0 0.4 34.7 4.0 6.0 3.5
SparQ 20.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 2.0 1.0 15.3 4.0 14.0 4.3
Hybrid 100.0 92.0 26.0 82.0 22.0 0.0 52.5 71.5 74.8 0.2 56.7 62.0 46.0 52.7
RocketKV 100.0 100.0 100.0 100.0 96.0 96.0 91.0 95.5 86.8 0.0 76.7 68.0 52.0 81.7
1024 Exact-TopK 100.0 100.0 100.0 100.0 100.0 96.0 98.5 98.5 93.6 0.2 74.7 74.0 54.0 83.8
DuoAttention 0.0 2.0 0.0 8.0 0.0 2.0 5.0 3.5 0.4 0.8 96.0 40.0 30.0 14.4
SnapKV 100.0 82.0 0.0 76.0 40.0 0.0 22.0 35.0 55.2 0.0 58.0 62.0 50.0 44.6
Quest 100.0 32.0 10.0 24.0 2.0 0.0 1.5 6.0 35.6 0.4 49.3 12.0 12.0 21.9
SparQ 48.0 0.0 0.0 2.0 0.0 0.0 0.5 0.0 12.4 0.8 18.7 20.0 24.0 9.7
Hybrid 100.0 100.0 100.0 100.0 70.0 2.0 92.0 98.0 84.4 0.2 70.0 72.0 54.0 72.5
RocketKV 100.0 100.0 100.0 100.0 96.0 100.0 92.0 98.0 87.2 0.0 78.7 72.0 54.0 82.9
2048 Exact-TopK 100.0 100.0 100.0 100.0 100.0 98.0 97.0 99.0 94.0 0.0 78.7 74.0 52.0 84.1
DuoAttention 0.0 2.0 0.0 8.0 0.0 2.0 5.0 3.5 0.4 0.8 94.0 44.0 28.0 14.4
SnapKV 100.0 82.0 4.0 98.0 52.0 0.0 50.5 72.0 82.0 0.4 60.0 70.0 50.0 55.5
Quest 100.0 78.0 70.0 88.0 48.0 0.0 19.5 39.5 76.8 0.0 63.3 38.0 38.0 50.7
SparQ 98.0 46.0 0.0 54.0 10.0 0.0 31.0 27.0 41.2 5.6 52.0 36.0 36.0 33.6
Hybrid 100.0 100.0 98.0 100.0 96.0 54.0 96.0 99.0 85.6 0.8 68.7 78.0 56.0 79.4
RocketKV 100.0 100.0 100.0 100.0 96.0 100.0 93.5 99.5 89.2 0.0 82.7 74.0 54.0 83.8
4096 Exact-TopK 100.0 100.0 100.0 100.0 100.0 100.0 96.5 99.0 95.6 0.0 80.0 76.0 52.0 84.5
DuoAttention 0.0 4.0 0.0 8.0 0.0 2.0 4.5 3.5 0.4 0.8 91.3 44.0 32.0 14.7
SnapKV 100.0 84.0 14.0 98.0 54.0 2.0 71.0 88.0 91.2 1.2 60.0 76.0 50.0 60.7
Quest 100.0 76.0 20.0 100.0 76.0 0.0 48.0 77.5 78.8 0.0 72.0 76.0 52.0 59.7
SparQ 100.0 100.0 100.0 100.0 94.0 32.0 94.0 96.5 86.4 2.4 73.3 70.0 52.0 77.0
Hybrid 100.0 100.0 100.0 100.0 98.0 96.0 96.5 98.0 90.8 0.2 70.0 76.0 56.0 83.2
RocketKV 100.0 100.0 100.0 100.0 98.0 100.0 98.0 99.5 92.4 0.0 84.0 74.0 56.0 84.8
Table 7: Results of RULER, SeqLen=96000 for Llama3.1-8B-Ins
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 100.0 100.0 100.0 94.0 88.0 98.5 100.0 66.8 0.0 74.7 78.0 50.0 80.8
256 Exact-TopK 100.0 100.0 100.0 100.0 84.0 14.0 95.5 99.0 67.2 0.4 56.0 82.0 50.0 72.9
DuoAttention 0.0 2.0 4.0 6.0 0.0 0.0 3.5 5.0 1.6 0.0 38.7 40.0 32.0 10.2
SnapKV 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 50.0 34.0 6.5
Quest 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.6 19.3 0.0 2.0 1.7
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Hybrid 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 4.8 0.0 11.3 4.0 4.0 1.9
RocketKV 100.0 98.0 56.0 94.0 66.0 0.0 47.5 54.5 53.6 0.0 51.3 80.0 46.0 57.5
512 Exact-TopK 100.0 100.0 100.0 100.0 84.0 32.0 96.5 100.0 67.6 0.4 59.3 80.0 50.0 74.6
DuoAttention 0.0 2.0 6.0 6.0 0.0 0.0 3.5 5.5 1.6 0.0 70.0 38.0 32.0 12.7
SnapKV 90.0 50.0 0.0 30.0 0.0 0.0 0.0 0.0 0.8 0.0 22.7 60.0 38.0 22.4
Quest 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 26.0 0.0 4.0 2.3
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Hybrid 94.0 46.0 0.0 18.0 0.0 0.0 4.0 7.5 42.0 0.6 28.0 20.0 16.0 21.2
RocketKV 100.0 100.0 96.0 100.0 80.0 6.0 89.0 95.0 58.4 0.0 56.7 78.0 44.0 69.5
1024 Exact-TopK 100.0 100.0 100.0 100.0 90.0 58.0 96.0 100.0 68.4 0.0 61.3 80.0 48.0 77.1
DuoAttention 0.0 2.0 6.0 6.0 0.0 0.0 3.5 5.5 1.6 0.0 70.7 40.0 32.0 12.9
SnapKV 100.0 78.0 0.0 62.0 14.0 0.0 5.0 12.0 32.8 0.0 33.3 58.0 40.0 33.5
Quest 94.0 0.0 0.0 10.0 0.0 0.0 2.5 2.5 0.4 0.2 58.0 2.0 8.0 13.7
SparQ 36.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.2 0.6 15.3 2.0 0.0 4.2
Hybrid 100.0 88.0 2.0 84.0 12.0 0.0 61.5 72.0 66.0 1.2 43.3 58.0 42.0 48.5
RocketKV 98.0 100.0 96.0 100.0 88.0 72.0 93.5 98.5 63.6 0.0 56.0 76.0 50.0 76.3
2048 Exact-TopK 100.0 100.0 100.0 100.0 92.0 82.0 95.5 100.0 64.4 0.0 66.7 80.0 48.0 79.1
DuoAttention 0.0 2.0 6.0 6.0 0.0 0.0 3.5 5.0 1.6 0.0 70.7 44.0 28.0 12.8
SnapKV 100.0 94.0 0.0 88.0 20.0 0.0 28.5 47.0 50.4 0.4 40.0 66.0 42.0 44.3
Quest 100.0 56.0 0.0 62.0 4.0 0.0 11.5 9.5 54.0 0.2 59.3 32.0 22.0 31.6
SparQ 12.0 0.0 0.0 0.0 0.0 0.0 0.0 2.0 3.6 0.2 32.7 8.0 14.0 5.6
Hybrid 100.0 96.0 74.0 98.0 54.0 0.0 89.0 96.5 77.2 0.4 52.0 78.0 48.0 66.4
RocketKV 100.0 100.0 100.0 100.0 88.0 84.0 96.0 99.5 65.6 0.0 56.7 76.0 48.0 78.0
4096 Exact-TopK 100.0 100.0 100.0 100.0 96.0 88.0 96.5 100.0 65.6 0.0 71.3 78.0 46.0 80.1
DuoAttention 0.0 2.0 6.0 8.0 0.0 0.0 3.5 5.0 1.2 0.0 72.0 44.0 32.0 13.4
SnapKV 100.0 100.0 0.0 96.0 36.0 0.0 50.5 75.5 62.8 0.2 49.3 74.0 42.0 52.8
Quest 98.0 100.0 0.0 94.0 36.0 2.0 40.5 63.0 45.2 0.0 58.7 36.0 36.0 46.9
SparQ 100.0 62.0 2.0 60.0 12.0 0.0 68.5 62.5 46.0 0.4 46.7 48.0 42.0 42.3
Hybrid 100.0 100.0 98.0 100.0 88.0 16.0 96.5 99.0 70.4 0.2 54.7 76.0 48.0 72.8
RocketKV 100.0 100.0 100.0 100.0 88.0 88.0 97.0 99.5 69.2 0.0 58.7 76.0 48.0 78.8
Table 8: Average results of LongBench and NIAH for Mistral-7B-Ins-v0.2
Token Budget
Method
Single QA
Multi. QA
Summ.
Few-shot
Synthetic
Code
LB Avg.
NIAH(%)
N/A Full-KV 36.4 29.6 28.1 66.7 44.9 55.6 45.4 99.0
256 Exact-TopK 35.5 28.6 28.2 67.6 42.6 58.2 45.3 98.7
DuoAttention 17.7 19.6 18.5 48.5 5.2 52.3 28.3 22.0
SnapKV 11.2 19.8 N/A 33.8 20.9 47.7 N/A 2.3
Quest 8.7 5.3 15.3 12.6 3.2 32.6 13.1 14.0
SparQ 1.6 0.9 8.6 2.5 6.4 26.0 7.0 10.0
Hybrid 29.7 21.9 25.5 58.5 25.0 50.9 37.0 56.7
RocketKV 34.4 26.8 24.6 65.6 38.5 57.8 42.9 97.7
512 Exact-TopK 35.8 29.2 28.2 67.1 42.9 57.4 45.3 99.3
DuoAttention 19.8 19.2 21.1 53.9 7.1 53.8 30.8 24.0
SnapKV 25.1 22.8 N/A 53.2 35.2 53.3 N/A 39.3
Quest 17.4 12.3 22.4 35.8 23.0 44.8 26.4 39.7
SparQ 5.1 1.7 13.0 8.2 2.3 35.3 10.6 11.7
Hybrid 34.2 27.4 27.2 65.9 38.3 57.2 43.5 85.7
RocketKV 35.3 27.9 25.8 66.6 41.2 57.9 44.1 98.7
1024 Exact-TopK 35.8 29.5 28.2 67.0 42.8 56.7 45.2 99.3
DuoAttention 21.2 19.4 22.4 59.1 8.1 55.0 32.6 30.0
SnapKV 29.2 25.3 20.7 59.6 42.1 56.2 39.9 53.3
Quest 25.4 18.6 25.5 56.2 36.3 53.3 36.9 79.3
SparQ 13.1 4.1 16.9 18.1 0.5 43.5 16.3 20.0
Hybrid 35.3 28.0 27.7 67.3 42.1 57.7 44.8 97.7
RocketKV 35.6 28.1 26.9 66.4 42.2 57.6 44.6 98.7
2048 Exact-TopK 36.2 29.7 28.2 66.6 42.9 56.8 45.2 99.3
DuoAttention 27.1 21.5 23.5 62.4 12.1 55.6 35.7 41.0
SnapKV 33.1 26.7 23.3 63.9 43.1 57.6 42.6 62.3
Quest 32.2 26.9 27.0 62.5 39.3 56.0 42.3 90.3
SparQ 22.0 10.3 20.1 34.1 4.0 50.5 24.5 27.0
Hybrid 36.2 29.0 28.2 67.2 42.2 57.2 45.2 99.0
RocketKV 35.9 29.1 27.6 66.9 42.5 57.2 45.0 99.3
4096 Exact-TopK 36.2 29.6 28.1 66.5 43.0 56.5 45.2 99.3
DuoAttention 32.1 25.3 25.6 65.4 26.7 56.4 40.6 62.0
SnapKV 34.1 28.0 25.2 64.6 42.8 57.9 43.6 76.7
Quest 35.1 29.1 27.9 66.3 42.7 56.3 44.7 96.0
SparQ 30.6 20.3 24.8 59.4 37.7 53.3 38.9 50.3
Hybrid 36.4 29.5 28.1 66.8 42.8 57.0 45.3 99.3
RocketKV 36.2 29.5 27.9 66.6 42.6 57.0 45.1 99.3
Table 9: Individual results of LongBench for Mistral-7B-Ins-v0.2
Token Budget
Method
narrativeqa
qasper
mul.fieldqa
hotpotqa
2wikimqa
musique
gov-report
qmsum
multi-news
trec
triviaqa
samsum
pass-ret.
lcc
repobench-p
passagecnt
N/A Full-KV 26.8 33.1 49.3 42.8 27.3 18.8 33.0 24.2 27.1 71.0 86.2 42.8 87.0 56.9 54.3 2.8
256 Exact-TopK 26.2 31.3 48.9 40.0 26.2 19.6 33.3 24.2 27.1 73.0 86.6 43.3 82.9 59.0 57.3 2.3
DuoAttention 13.0 11.7 28.5 29.5 20.1 9.3 17.6 19.4 18.6 26.0 81.8 37.7 7.1 50.4 54.1 3.2
SnapKV 0.9 1.9 30.7 26.4 21.3 11.6 N/A N/A N/A 23.0 77.9 0.5 38.0 46.0 49.5 3.8
Quest 2.1 9.1 14.9 6.0 7.8 2.1 13.0 10.8 22.2 14.0 14.1 9.6 5.9 39.6 25.6 0.4
SparQ 0.3 1.7 2.9 0.9 1.5 0.1 5.3 6.7 13.8 0.0 7.1 0.4 12.7 30.0 22.0 0.2
Hybrid 17.1 28.1 43.9 30.2 23.6 11.8 27.3 22.4 26.8 63.0 78.8 33.6 46.9 52.1 49.7 3.2
RocketKV 25.6 28.8 48.9 37.2 24.8 18.4 26.0 23.6 24.3 68.5 86.3 42.0 73.7 57.2 58.5 3.3
512 Exact-TopK 27.2 31.6 48.6 40.5 27.5 19.7 33.2 24.2 27.1 72.0 86.6 42.8 83.1 58.3 56.5 2.6
DuoAttention 14.7 13.8 30.8 27.2 20.5 10.0 20.6 19.9 22.8 40.0 81.8 39.8 11.7 53.5 54.0 2.5
SnapKV 18.1 15.8 41.2 32.7 21.4 14.4 N/A N/A N/A 38.0 83.2 38.4 66.7 52.4 54.3 3.6
Quest 6.0 17.0 29.1 18.1 13.8 5.1 22.8 18.9 25.5 35.0 50.7 21.8 43.1 48.3 41.2 2.9
SparQ 0.1 4.1 10.9 1.8 2.8 0.5 6.8 11.9 20.3 4.5 16.2 3.9 4.4 45.2 25.4 0.2
Hybrid 23.2 31.2 48.4 38.6 26.6 17.0 31.0 23.3 27.2 71.5 85.0 41.2 74.7 57.1 57.3 1.9
RocketKV 26.5 30.6 48.9 38.7 25.9 19.0 28.0 24.3 25.1 70.5 86.1 43.3 79.2 58.2 57.6 3.3
1024 Exact-TopK 27.2 31.3 48.8 41.7 27.4 19.3 33.4 24.1 27.2 72.0 86.4 42.5 82.8 57.7 55.7 2.9
DuoAttention 15.7 15.7 32.3 29.4 19.3 9.4 22.0 20.4 24.8 53.0 83.8 40.5 13.2 55.1 54.9 3.1
SnapKV 20.0 22.4 45.2 36.3 23.1 16.6 18.5 22.1 21.6 52.5 85.6 40.6 80.9 55.3 57.2 3.3
Quest 10.1 26.7 39.6 26.5 17.9 11.4 27.5 22.9 26.3 66.0 70.5 32.3 69.8 55.7 50.9 2.8
SparQ 0.2 14.5 24.5 5.8 6.0 0.5 9.3 16.3 25.1 13.0 31.1 10.1 1.0 54.3 32.8 0.0
Hybrid 26.1 31.6 48.2 38.4 26.8 18.8 32.1 23.5 27.3 73.0 86.5 42.5 81.8 58.8 56.7 2.5
RocketKV 26.4 32.1 48.5 39.9 25.9 18.5 29.7 24.6 26.4 71.0 86.1 42.1 82.1 58.4 56.9 2.4
2048 Exact-TopK 27.3 32.8 48.6 41.8 27.9 19.4 33.3 24.0 27.4 71.0 86.4 42.4 82.8 57.9 55.6 3.0
DuoAttention 17.2 24.5 39.5 32.9 21.8 9.9 23.5 20.6 26.6 61.5 84.3 41.3 20.7 56.1 55.0 3.4
SnapKV 24.8 26.9 47.7 38.1 23.6 18.4 22.4 23.5 24.1 65.5 85.7 40.7 82.9 57.6 57.7 3.3
Quest 18.3 31.5 46.7 36.9 25.5 18.3 31.0 23.1 27.0 70.5 79.6 37.4 76.3 58.0 54.1 2.3
SparQ 2.9 29.3 33.8 9.0 21.4 0.6 16.7 16.9 26.6 35.5 47.6 19.2 6.9 57.2 43.8 1.2
Hybrid 27.2 32.2 49.1 40.8 27.0 19.2 33.2 24.1 27.4 72.0 86.8 42.9 81.3 57.7 56.8 3.1
RocketKV 27.7 32.0 48.1 41.2 26.7 19.5 31.4 24.4 26.9 71.5 86.4 42.8 82.7 58.1 56.2 2.4
4096 Exact-TopK 27.7 32.7 48.2 41.5 27.9 19.4 33.1 24.1 27.2 71.0 86.2 42.4 83.3 57.8 55.3 2.8
DuoAttention 19.1 31.6 45.6 36.6 26.8 12.4 27.7 21.8 27.2 70.0 85.0 41.2 51.2 56.8 56.1 2.2
SnapKV 24.9 29.7 47.7 39.6 25.0 19.5 25.6 23.9 26.1 66.5 86.0 41.3 82.8 58.7 57.0 2.8
Quest 24.6 32.1 48.7 40.4 27.1 19.7 32.8 23.5 27.2 72.5 86.3 40.3 82.7 57.7 54.9 2.8
SparQ 12.3 31.7 47.9 26.1 25.8 9.0 26.5 20.7 27.1 71.0 69.6 37.5 71.5 57.4 49.1 3.9
Hybrid 27.8 32.2 49.0 42.0 27.4 19.2 33.0 23.9 27.2 71.5 86.2 42.7 83.3 57.9 56.2 2.3
RocketKV 27.7 32.1 48.7 41.2 27.8 19.6 32.2 24.6 27.0 71.5 86.2 42.0 82.2 57.8 56.2 3.1
Table 10: Results of RULER, SeqLen=8000 for Mistral-7B-Ins-v0.2
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 100.0 100.0 98.0 100.0 82.0 79.5 94.0 97.6 90.6 71.3 82.0 64.0 89.2
256 Exact-TopK 100.0 100.0 100.0 98.0 100.0 84.0 80.5 98.0 96.0 88.0 68.7 84.0 66.0 89.5
DuoAttention 4.0 0.0 2.0 2.0 0.0 0.0 2.0 2.0 2.4 26.8 48.7 32.0 40.0 12.5
SnapKV 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 3.6 1.3 48.0 50.0 7.9
Quest 92.0 32.0 0.0 24.0 6.0 0.0 3.0 6.0 15.2 19.6 38.0 22.0 28.0 22.0
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 2.2 9.3 2.0 10.0 1.8
Hybrid 98.0 82.0 50.0 84.0 74.0 12.0 42.5 60.5 88.4 30.4 73.3 84.0 66.0 65.0
RocketKV 100.0 100.0 98.0 98.0 100.0 72.0 82.0 95.5 93.6 47.8 74.0 88.0 62.0 85.5
512 Exact-TopK 100.0 100.0 100.0 98.0 100.0 82.0 80.5 98.0 96.0 89.6 72.7 82.0 66.0 89.6
DuoAttention 8.0 12.0 2.0 4.0 4.0 8.0 9.5 8.0 6.0 2.4 69.3 26.0 42.0 15.5
SnapKV 18.0 36.0 0.0 10.0 34.0 0.0 0.0 0.0 11.6 6.6 52.7 76.0 54.0 23.0
Quest 94.0 76.0 2.0 92.0 52.0 0.0 53.5 59.5 76.4 29.8 65.3 64.0 54.0 55.3
SparQ 0.0 0.0 0.0 0.0 0.0 6.0 0.0 0.0 0.0 2.6 67.3 0.0 32.0 8.3
Hybrid 100.0 98.0 96.0 98.0 96.0 80.0 75.5 93.0 95.2 58.0 71.3 84.0 66.0 85.5
RocketKV 100.0 100.0 100.0 98.0 100.0 80.0 87.5 97.5 96.8 59.6 74.7 84.0 66.0 88.0
1024 Exact-TopK 100.0 100.0 100.0 98.0 100.0 82.0 80.0 96.0 96.8 90.8 73.3 80.0 66.0 89.5
DuoAttention 12.0 12.0 2.0 4.0 14.0 10.0 9.5 8.0 9.2 14.2 72.0 42.0 48.0 19.8
SnapKV 92.0 66.0 0.0 54.0 64.0 0.0 3.5 7.5 79.6 12.0 73.3 78.0 58.0 45.2
Quest 92.0 84.0 26.0 94.0 76.0 22.0 74.5 78.0 89.6 51.8 70.0 84.0 64.0 69.7
SparQ 0.0 62.0 6.0 72.0 0.0 72.0 70.0 61.0 0.0 8.6 68.0 30.0 44.0 38.0
Hybrid 100.0 100.0 100.0 98.0 100.0 84.0 79.5 98.0 96.4 80.0 68.7 82.0 66.0 88.7
RocketKV 100.0 100.0 98.0 98.0 100.0 78.0 89.0 98.5 96.8 72.4 72.0 86.0 66.0 88.8
2048 Exact-TopK 100.0 100.0 100.0 98.0 100.0 82.0 79.5 95.5 97.6 91.2 74.0 80.0 66.0 89.5
DuoAttention 38.0 40.0 30.0 34.0 60.0 24.0 15.0 22.0 24.0 48.0 70.0 62.0 62.0 40.7
SnapKV 100.0 84.0 0.0 82.0 88.0 0.0 8.5 26.5 96.4 25.0 75.3 78.0 66.0 56.1
Quest 100.0 96.0 98.0 94.0 98.0 80.0 80.5 95.5 97.2 74.2 70.7 82.0 66.0 87.1
SparQ 96.0 90.0 100.0 96.0 100.0 84.0 78.5 95.0 81.6 90.0 70.7 82.0 66.0 86.9
Hybrid 100.0 100.0 100.0 98.0 100.0 82.0 80.5 96.0 96.4 88.4 69.3 80.0 66.0 89.0
RocketKV 100.0 100.0 100.0 98.0 100.0 78.0 90.5 98.0 98.4 82.6 74.0 82.0 68.0 90.0
4096 Exact-TopK 100.0 100.0 100.0 98.0 100.0 82.0 79.5 94.5 97.6 90.8 73.3 82.0 64.0 89.4
DuoAttention 100.0 98.0 100.0 96.0 100.0 84.0 71.0 90.5 96.0 85.4 70.0 86.0 66.0 87.9
SnapKV 100.0 96.0 2.0 94.0 90.0 6.0 24.0 51.5 98.4 52.2 73.3 80.0 66.0 64.1
Quest 100.0 100.0 100.0 98.0 100.0 84.0 80.0 96.0 96.8 87.8 71.3 82.0 66.0 89.4
SparQ 100.0 100.0 100.0 98.0 100.0 82.0 79.0 95.5 97.6 89.8 72.7 82.0 64.0 89.3
Hybrid 100.0 100.0 100.0 98.0 100.0 82.0 79.5 96.0 97.6 90.6 74.0 80.0 64.0 89.4
RocketKV 100.0 100.0 94.0 98.0 100.0 82.0 91.0 98.0 98.8 87.2 74.0 84.0 66.0 90.2
Table 11: Results of RULER, SeqLen=16000 for Mistral-7B-Ins-v0.2
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 100.0 100.0 86.0 90.0 74.0 84.0 85.5 87.6 74.2 86.7 78.0 54.0 84.6
256 Exact-TopK 100.0 100.0 100.0 84.0 92.0 82.0 87.0 80.0 85.2 70.6 76.7 80.0 58.0 84.3
DuoAttention 2.0 0.0 2.0 2.0 0.0 0.0 2.5 3.5 0.8 12.6 48.7 30.0 36.0 10.8
SnapKV 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 2.6 0.7 38.0 40.0 6.3
Quest 0.0 2.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 3.6 18.7 8.0 12.0 3.4
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.4 0.0 2.0 0.0 0.2
Hybrid 96.0 52.0 2.0 30.0 24.0 0.0 16.0 10.0 72.0 16.0 70.7 74.0 54.0 39.7
RocketKV 100.0 100.0 98.0 86.0 92.0 42.0 79.5 91.5 85.6 37.0 79.3 76.0 52.0 78.4
512 Exact-TopK 100.0 100.0 100.0 84.0 92.0 80.0 85.0 84.0 86.0 72.6 79.3 80.0 58.0 84.7
DuoAttention 2.0 2.0 4.0 2.0 0.0 0.0 6.5 4.5 2.8 0.4 80.7 32.0 32.0 13.0
SnapKV 26.0 18.0 0.0 2.0 12.0 0.0 0.0 0.0 10.4 2.2 50.0 62.0 46.0 17.6
Quest 100.0 72.0 4.0 56.0 12.0 0.0 18.5 16.5 41.2 13.4 62.7 22.0 36.0 34.9
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.2 2.7 0.0 6.0 0.8
Hybrid 100.0 94.0 86.0 80.0 76.0 18.0 47.5 59.0 81.6 34.4 75.3 80.0 58.0 68.4
RocketKV 100.0 100.0 98.0 82.0 86.0 44.0 87.5 95.0 84.8 48.0 81.3 80.0 52.0 79.9
1024 Exact-TopK 100.0 100.0 100.0 84.0 92.0 78.0 83.0 85.5 88.4 74.4 84.0 78.0 58.0 85.0
DuoAttention 6.0 8.0 4.0 2.0 0.0 0.0 8.0 5.0 4.4 0.2 80.7 38.0 46.0 15.6
SnapKV 90.0 62.0 0.0 36.0 36.0 0.0 3.0 3.5 73.2 8.4 73.3 64.0 48.0 38.3
Quest 84.0 94.0 10.0 82.0 44.0 0.0 61.5 54.5 73.6 28.8 70.0 68.0 56.0 55.9
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 6.6 60.0 0.0 2.0 5.3
Hybrid 100.0 100.0 98.0 86.0 90.0 72.0 76.5 77.5 82.4 57.8 78.0 78.0 56.0 80.9
RocketKV 100.0 100.0 98.0 82.0 88.0 56.0 93.0 97.0 88.4 56.8 83.3 76.0 52.0 82.3
2048 Exact-TopK 100.0 100.0 100.0 84.0 92.0 76.0 83.0 85.5 88.8 74.2 86.7 80.0 58.0 85.2
DuoAttention 10.0 8.0 6.0 6.0 6.0 0.0 8.0 6.0 4.8 2.2 79.3 40.0 48.0 17.3
SnapKV 98.0 76.0 0.0 54.0 44.0 0.0 9.5 17.0 85.6 22.0 81.3 72.0 50.0 46.9
Quest 92.0 88.0 44.0 76.0 72.0 10.0 68.0 73.0 84.0 51.6 74.7 80.0 58.0 67.0
SparQ 0.0 6.0 0.0 16.0 0.0 4.0 22.0 6.0 0.4 12.8 79.3 4.0 20.0 13.1
Hybrid 100.0 100.0 100.0 84.0 92.0 76.0 82.5 82.0 85.2 70.8 78.0 76.0 58.0 83.4
RocketKV 100.0 100.0 100.0 82.0 88.0 66.0 93.0 96.5 87.2 65.8 84.7 76.0 56.0 84.2
4096 Exact-TopK 100.0 100.0 100.0 84.0 92.0 74.0 83.0 86.0 88.0 75.2 86.7 80.0 56.0 85.0
DuoAttention 48.0 46.0 40.0 32.0 48.0 14.0 24.0 20.0 22.8 16.6 80.7 62.0 58.0 39.4
SnapKV 100.0 88.0 2.0 74.0 58.0 0.0 20.0 39.0 92.0 36.4 81.3 74.0 52.0 55.1
Quest 100.0 100.0 96.0 86.0 92.0 82.0 82.5 83.0 86.0 65.8 80.7 80.0 58.0 84.0
SparQ 82.0 86.0 84.0 82.0 80.0 78.0 67.5 64.5 31.2 73.2 83.3 76.0 52.0 72.3
Hybrid 100.0 100.0 100.0 84.0 92.0 78.0 84.0 86.5 84.4 74.6 81.3 78.0 58.0 84.7
RocketKV 100.0 100.0 100.0 82.0 88.0 70.0 95.5 98.0 88.8 70.0 86.0 74.0 56.0 85.3
Table 12: Results of RULER, SeqLen=24000 for Mistral-7B-Ins-v0.2
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 100.0 92.0 86.0 94.0 78.0 77.5 83.5 93.2 40.6 86.7 68.0 58.0 81.3
256 Exact-TopK 100.0 98.0 88.0 90.0 92.0 78.0 78.5 79.0 92.8 62.4 80.0 70.0 58.0 82.1
DuoAttention 2.0 0.0 2.0 6.0 2.0 0.0 1.5 3.0 0.8 1.2 56.7 24.0 34.0 10.2
SnapKV 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.4 0.0 38.0 38.0 6.0
Quest 0.0 2.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.8 15.3 2.0 2.0 1.8
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 6.0 0.0 0.5
Hybrid 90.0 34.0 0.0 24.0 0.0 0.0 8.0 7.5 73.6 6.4 56.0 52.0 44.0 30.4
RocketKV 100.0 98.0 90.0 78.0 88.0 36.0 73.0 79.5 87.2 18.8 67.3 76.0 52.0 72.6
512 Exact-TopK 100.0 100.0 90.0 88.0 92.0 80.0 77.0 81.0 93.2 59.0 86.0 70.0 58.0 82.6
DuoAttention 4.0 4.0 4.0 6.0 6.0 2.0 4.5 4.5 1.6 0.4 68.7 26.0 28.0 12.3
SnapKV 42.0 20.0 0.0 6.0 10.0 0.0 0.0 0.0 5.2 1.6 38.0 50.0 42.0 16.5
Quest 100.0 12.0 32.0 8.0 6.0 0.0 2.0 7.0 7.2 3.0 56.7 8.0 16.0 19.8
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.6 0.0 2.0 0.0 0.2
Hybrid 100.0 74.0 48.0 54.0 42.0 0.0 33.0 33.5 86.0 18.0 66.7 68.0 52.0 51.9
RocketKV 100.0 100.0 96.0 84.0 96.0 48.0 80.5 91.5 95.2 32.4 74.0 70.0 54.0 78.6
1024 Exact-TopK 100.0 100.0 90.0 86.0 92.0 80.0 78.0 84.0 93.6 55.4 86.7 70.0 58.0 82.6
DuoAttention 4.0 4.0 4.0 6.0 8.0 2.0 4.5 4.5 2.0 0.6 70.0 28.0 32.0 13.0
SnapKV 96.0 58.0 0.0 24.0 30.0 0.0 2.5 4.0 68.0 4.2 65.3 60.0 46.0 35.2
Quest 100.0 82.0 2.0 78.0 20.0 0.0 32.0 34.5 45.2 9.6 66.0 36.0 32.0 41.3
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 2.2 1.3 0.0 4.0 0.6
Hybrid 100.0 98.0 86.0 80.0 72.0 36.0 62.0 71.5 93.2 36.4 72.7 72.0 56.0 72.0
RocketKV 100.0 100.0 98.0 82.0 94.0 50.0 83.0 93.5 96.8 38.2 78.0 68.0 54.0 79.7
2048 Exact-TopK 100.0 100.0 90.0 86.0 94.0 80.0 78.0 83.0 93.6 50.2 88.7 70.0 58.0 82.4
DuoAttention 4.0 4.0 4.0 6.0 10.0 2.0 4.5 4.5 2.4 0.4 69.3 26.0 30.0 12.9
SnapKV 96.0 80.0 0.0 52.0 40.0 0.0 7.0 17.5 96.0 11.0 70.7 64.0 46.0 44.6
Quest 96.0 88.0 2.0 80.0 64.0 0.0 65.0 66.0 88.0 18.0 75.3 62.0 50.0 58.0
SparQ 0.0 0.0 0.0 0.0 0.0 34.0 0.0 0.0 0.0 27.2 80.0 2.0 4.0 11.3
Hybrid 100.0 98.0 88.0 88.0 90.0 68.0 73.5 80.5 94.8 53.6 74.7 68.0 60.0 79.8
RocketKV 100.0 100.0 98.0 84.0 94.0 64.0 87.5 93.5 98.4 44.8 82.7 60.0 54.0 81.6
4096 Exact-TopK 100.0 100.0 90.0 86.0 94.0 78.0 78.0 84.0 93.6 41.2 88.7 70.0 58.0 81.7
DuoAttention 24.0 18.0 10.0 18.0 20.0 2.0 6.0 6.0 6.4 0.8 71.3 52.0 44.0 21.4
SnapKV 100.0 84.0 0.0 60.0 54.0 2.0 19.5 37.5 96.4 21.8 75.3 68.0 46.0 51.1
Quest 90.0 86.0 72.0 72.0 72.0 38.0 71.5 75.0 90.4 43.4 80.0 70.0 56.0 70.5
SparQ 0.0 6.0 0.0 4.0 0.0 8.0 2.0 10.5 0.0 6.2 86.0 0.0 18.0 10.8
Hybrid 100.0 98.0 90.0 88.0 92.0 78.0 73.5 83.0 94.8 62.0 82.7 70.0 58.0 82.3
RocketKV 100.0 100.0 98.0 82.0 92.0 70.0 87.5 94.0 97.2 47.8 85.3 60.0 54.0 82.1
Table 13: Results of RULER, SeqLen=32000 for Mistral-7B-Ins-v0.2
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 98.0 86.0 82.0 88.0 68.0 83.0 81.0 90.0 29.6 94.7 70.0 48.0 78.3
256 Exact-TopK 100.0 92.0 68.0 78.0 80.0 60.0 78.0 58.5 91.2 40.4 86.0 66.0 48.0 72.8
DuoAttention 0.0 2.0 4.0 4.0 0.0 0.0 1.0 3.0 0.0 2.0 34.0 24.0 30.0 8.0
SnapKV 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.6 0.0 22.0 36.0 4.5
Quest 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.6 22.0 0.0 2.0 1.9
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 2.0 0.0 0.2
Hybrid 76.0 14.0 0.0 4.0 2.0 0.0 6.0 3.5 55.6 2.2 57.3 26.0 32.0 21.4
RocketKV 98.0 86.0 50.0 68.0 66.0 16.0 48.5 44.5 82.8 11.2 78.7 70.0 42.0 58.6
512 Exact-TopK 100.0 94.0 68.0 78.0 86.0 64.0 83.0 60.5 93.2 37.4 87.3 66.0 46.0 74.1
DuoAttention 2.0 6.0 6.0 4.0 0.0 2.0 3.5 4.0 2.8 0.8 75.3 24.0 30.0 12.3
SnapKV 56.0 6.0 0.0 6.0 6.0 0.0 0.0 0.0 2.8 1.0 30.0 46.0 38.0 14.8
Quest 88.0 2.0 4.0 2.0 0.0 0.0 1.5 1.0 0.0 4.0 64.0 8.0 16.0 14.7
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.4 0.0 0.0 2.0 0.2
Hybrid 96.0 28.0 4.0 28.0 26.0 0.0 16.5 12.5 77.2 9.2 77.3 72.0 44.0 37.7
RocketKV 100.0 100.0 80.0 80.0 90.0 38.0 84.0 81.5 92.0 16.8 83.3 66.0 48.0 73.8
1024 Exact-TopK 100.0 96.0 72.0 80.0 90.0 64.0 81.5 66.0 92.4 37.0 91.3 68.0 50.0 76.0
DuoAttention 2.0 6.0 6.0 4.0 0.0 2.0 3.5 4.5 3.6 0.6 80.7 20.0 32.0 12.7
SnapKV 100.0 46.0 0.0 20.0 14.0 0.0 1.5 2.5 44.0 3.0 70.0 58.0 40.0 30.7
Quest 94.0 74.0 2.0 44.0 18.0 0.0 15.5 19.0 69.6 5.6 77.3 28.0 32.0 36.8
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.6 0.0 0.0 8.0 0.7
Hybrid 98.0 70.0 56.0 56.0 50.0 10.0 46.5 38.0 85.2 24.8 78.7 74.0 48.0 56.6
RocketKV 100.0 100.0 84.0 80.0 90.0 28.0 86.5 87.0 94.0 23.0 84.7 68.0 48.0 74.9
2048 Exact-TopK 100.0 98.0 76.0 82.0 90.0 64.0 84.0 73.5 90.8 34.0 92.7 68.0 48.0 77.0
DuoAttention 2.0 6.0 6.0 4.0 0.0 2.0 3.5 5.0 3.6 0.8 81.3 28.0 30.0 13.2
SnapKV 100.0 66.0 0.0 44.0 20.0 0.0 4.5 10.0 85.2 7.2 78.0 62.0 42.0 39.9
Quest 98.0 94.0 2.0 70.0 32.0 0.0 49.5 46.0 91.2 7.0 84.7 52.0 40.0 51.3
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 4.4 67.3 0.0 4.0 5.8
Hybrid 100.0 96.0 76.0 72.0 72.0 40.0 71.5 59.5 90.8 44.2 82.0 68.0 50.0 70.9
RocketKV 100.0 100.0 86.0 86.0 90.0 48.0 87.0 93.5 96.0 25.8 86.7 68.0 48.0 78.1
4096 Exact-TopK 100.0 98.0 76.0 82.0 92.0 64.0 82.5 77.0 90.4 34.8 93.3 68.0 48.0 77.4
DuoAttention 10.0 8.0 8.0 6.0 6.0 2.0 4.0 4.5 4.8 1.8 84.0 34.0 32.0 15.8
SnapKV 100.0 70.0 0.0 58.0 30.0 0.0 12.5 22.0 92.0 15.8 80.0 64.0 46.0 45.4
Quest 92.0 78.0 32.0 68.0 64.0 8.0 67.0 59.0 90.0 24.8 89.3 68.0 46.0 60.5
SparQ 0.0 2.0 0.0 0.0 0.0 0.0 0.5 0.0 0.0 9.4 88.0 0.0 4.0 8.0
Hybrid 100.0 98.0 76.0 78.0 84.0 60.0 74.0 68.5 94.4 43.4 85.3 66.0 48.0 75.0
RocketKV 100.0 100.0 86.0 88.0 90.0 56.0 88.0 93.0 96.4 30.0 88.7 66.0 48.0 79.2
Table 14: Average results of LongBench and NIAH for LongChat-7B-v1.5
Token Budget
Method
Single QA
Multi. QA
Summ.
Few-shot
Synthetic
Code
LB Avg.
NIAH(%)
N/A Full-KV 31.1 24.1 26.7 57.7 15.2 57.0 37.5 96.3
256 Exact-TopK 30.5 23.5 26.5 53.9 14.8 58.3 36.6 96.7
DuoAttention 14.7 15.8 14.4 29.8 2.4 48.5 21.7 20.7
SnapKV 24.4 21.8 N/A 42.7 15.0 52.2 N/A 16.7
Quest 9.2 7.2 7.8 11.4 0.9 34.4 11.8 15.3
SparQ 18.9 11.2 15.2 16.8 1.5 32.4 16.9 21.0
Hybrid 28.2 23.7 25.1 49.4 13.0 54.4 34.1 58.0
RocketKV 29.6 22.5 23.1 50.9 7.0 57.9 33.8 93.7
512 Exact-TopK 30.8 23.9 26.7 54.4 14.5 57.6 36.8 96.3
DuoAttention 18.1 20.4 19.5 37.7 3.2 49.4 26.1 23.7
SnapKV 27.8 23.5 N/A 50.0 13.5 55.0 N/A 58.3
Quest 20.7 15.6 16.8 32.5 7.6 46.4 24.2 48.0
SparQ 26.5 21.4 23.6 41.5 13.5 49.1 30.9 38.0
Hybrid 30.6 24.8 26.5 52.3 14.2 57.2 36.3 90.7
RocketKV 30.3 23.1 24.0 52.6 8.8 57.7 34.9 95.7
1024 Exact-TopK 30.7 23.9 26.8 54.3 14.2 57.4 36.7 96.3
DuoAttention 20.8 20.7 21.0 42.0 4.0 53.6 28.5 28.0
SnapKV 28.7 24.0 22.2 52.6 14.5 56.2 34.9 76.0
Quest 28.2 23.0 23.6 47.9 17.2 53.3 33.8 77.0
SparQ 29.7 22.9 26.4 53.3 17.2 55.9 36.1 76.7
Hybrid 30.7 24.2 27.0 53.8 14.0 57.9 36.7 95.7
RocketKV 30.4 23.0 25.2 54.0 11.5 57.4 35.7 96.3
2048 Exact-TopK 30.7 24.0 26.8 54.4 15.0 57.3 36.8 96.3
DuoAttention 26.0 20.2 23.1 46.3 3.8 55.0 30.9 42.3
SnapKV 29.9 24.1 24.4 54.2 14.5 56.6 36.0 85.0
Quest 30.0 24.2 26.5 53.3 15.5 56.3 36.3 89.7
SparQ 30.6 24.1 26.9 54.6 14.5 56.8 36.7 96.0
Hybrid 30.7 24.1 26.8 54.6 14.5 57.3 36.8 96.3
RocketKV 31.1 23.8 26.2 54.3 13.2 57.1 36.4 96.3
4096 Exact-TopK 31.0 24.0 26.6 54.4 15.5 57.2 36.9 96.3
DuoAttention 28.6 22.4 25.4 50.5 10.5 55.2 34.1 67.3
SnapKV 30.9 24.2 25.9 55.0 14.8 57.0 36.8 89.0
Quest 30.7 24.2 27.0 54.3 14.2 57.3 36.8 96.3
SparQ 30.8 23.9 26.7 54.6 15.0 57.2 36.8 94.3
Hybrid 31.0 23.9 26.8 54.4 14.2 57.3 36.8 96.3
RocketKV 31.0 24.2 26.5 54.7 14.2 57.2 36.8 96.0
Table 15: Individual results of LongBench for LongChat-7B-v1.5
Token Budget
Method
narrativeqa
qasper
mul.fieldqa
hotpotqa
2wikimqa
musique
gov-report
qmsum
multi-news
trec
triviaqa
samsum
pass-ret.
lcc
repobench-p
passagecnt
N/A Full-KV 20.8 29.4 43.1 33.0 24.4 14.7 30.8 22.8 26.7 66.5 84.0 22.5 30.5 54.7 59.2 0.0
256 Exact-TopK 19.6 30.2 41.6 33.7 23.5 13.5 30.6 22.5 26.3 65.5 83.5 12.7 29.5 57.7 58.9 0.0
DuoAttention 12.4 10.1 21.6 18.6 20.5 8.1 12.6 18.7 11.8 13.5 65.9 9.8 4.8 48.6 48.3 0.0
SnapKV 15.8 20.2 37.3 29.9 23.4 12.2 N/A N/A N/A 39.5 83.0 5.5 28.5 51.7 52.6 1.5
Quest 1.7 13.8 12.0 8.3 12.4 0.9 2.9 4.1 16.5 18.0 14.7 1.5 1.1 43.1 25.6 0.8
SparQ 2.5 27.3 26.8 12.9 19.4 1.2 10.4 9.8 25.3 22.5 25.4 2.5 2.5 44.8 20.0 0.5
Hybrid 14.1 29.8 40.7 35.0 23.8 12.3 27.3 21.3 26.7 62.0 78.3 7.8 24.1 57.8 50.9 2.0
RocketKV 18.8 28.6 41.3 31.0 23.4 13.3 24.6 22.1 22.7 62.0 82.5 8.1 13.5 57.4 58.4 0.5
512 Exact-TopK 20.1 30.5 41.9 33.9 23.8 14.0 30.6 22.9 26.6 66.0 84.4 12.9 29.0 56.4 58.8 0.0
DuoAttention 13.1 15.9 25.4 28.0 24.1 9.2 19.4 19.3 19.7 31.5 72.2 9.5 5.5 50.6 48.3 1.0
SnapKV 17.5 25.8 40.2 33.4 23.6 13.6 N/A N/A N/A 55.5 84.8 9.7 27.0 53.5 56.5 0.0
Quest 5.0 26.8 30.2 21.3 19.8 5.8 11.4 13.9 25.2 42.5 51.4 3.5 13.8 54.0 38.8 1.5
SparQ 10.3 29.6 39.5 30.6 24.1 9.6 23.9 20.5 26.5 62.0 58.0 4.6 26.1 54.6 43.5 1.0
Hybrid 19.1 29.8 42.8 35.1 24.3 15.1 30.8 22.4 26.4 64.5 83.6 8.8 28.0 57.3 57.2 0.5
RocketKV 20.3 28.9 41.8 32.3 23.7 13.4 26.3 22.3 23.3 65.0 83.2 9.4 17.5 57.2 58.1 0.0
1024 Exact-TopK 20.5 29.7 42.0 33.5 24.2 14.2 30.9 23.2 26.4 65.5 84.4 13.1 28.5 55.7 59.1 0.0
DuoAttention 13.5 18.1 30.7 28.2 24.8 9.2 20.2 19.9 22.9 43.5 72.8 9.7 6.5 56.6 50.6 1.5
SnapKV 19.3 26.9 40.1 35.5 23.2 13.3 21.4 22.0 23.1 62.5 83.8 11.4 29.0 53.9 58.4 0.0
Quest 13.2 29.5 42.0 33.1 24.4 11.5 23.6 21.4 25.9 63.0 76.5 4.1 32.0 55.9 50.8 2.5
SparQ 17.8 29.6 41.8 32.6 23.3 12.7 30.3 22.7 26.3 66.5 83.2 10.3 33.0 55.6 56.2 1.5
Hybrid 20.4 30.5 41.2 34.2 23.7 14.9 31.4 23.0 26.5 65.5 84.6 11.4 28.0 57.2 58.6 0.0
RocketKV 20.5 28.6 42.1 32.6 23.5 12.9 28.0 22.6 25.1 67.5 83.2 11.3 22.5 56.4 58.5 0.5
2048 Exact-TopK 20.3 29.5 42.4 33.2 24.7 14.2 31.1 22.9 26.5 66.0 84.2 12.9 30.0 55.3 59.4 0.0
DuoAttention 13.9 26.6 37.4 27.8 25.4 7.4 23.3 19.9 26.1 55.5 74.4 9.1 6.5 56.5 53.5 1.0
SnapKV 20.3 28.2 41.2 33.7 24.3 14.5 25.2 22.6 25.4 65.0 84.8 12.8 29.0 54.5 58.7 0.0
Quest 18.4 30.4 41.3 34.7 23.9 13.9 30.2 22.7 26.5 67.5 83.5 8.9 30.0 56.2 56.5 1.0
SparQ 20.3 29.5 42.0 33.3 24.7 14.4 31.3 22.9 26.5 66.0 84.5 13.3 29.0 55.1 58.4 0.0
Hybrid 20.3 29.9 42.0 34.3 24.2 13.8 31.2 22.6 26.4 67.0 84.4 12.4 29.0 55.5 59.1 0.0
RocketKV 20.6 29.1 43.5 32.6 24.5 14.2 29.6 22.8 26.1 67.5 83.8 11.6 26.5 55.1 59.1 0.0
4096 Exact-TopK 21.0 29.4 42.6 33.0 24.3 14.6 30.7 22.6 26.5 66.5 84.0 12.8 31.0 55.1 59.4 0.0
DuoAttention 14.2 29.6 42.1 32.2 23.9 11.1 28.6 21.1 26.5 64.0 78.3 9.1 20.5 55.7 54.6 0.5
SnapKV 20.6 29.4 42.7 34.1 24.2 14.4 28.3 23.0 26.4 67.5 84.2 13.3 29.5 55.0 59.1 0.0
Quest 20.5 29.8 41.9 33.9 24.8 14.0 31.4 22.8 26.7 67.0 84.3 11.6 28.5 55.4 59.2 0.0
SparQ 21.0 29.0 42.5 33.2 24.2 14.3 30.7 22.8 26.6 66.5 84.2 13.1 30.0 55.1 59.4 0.0
Hybrid 20.8 29.2 43.2 33.5 24.2 14.1 31.0 22.9 26.5 66.0 84.2 12.9 28.5 55.3 59.3 0.0
RocketKV 20.1 29.8 43.2 33.2 25.1 14.4 30.2 22.7 26.6 67.0 84.3 12.8 28.5 54.9 59.4 0.0
Table 16: Results of RULER, SeqLen=8000 for LongChat-7B-v1.5
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 100.0 96.0 98.0 100.0 74.0 96.0 69.0 94.0 33.8 80.7 70.0 58.0 82.3
256 Exact-TopK 100.0 96.0 96.0 94.0 100.0 84.0 88.0 60.0 90.8 39.4 76.7 66.0 58.0 80.7
DuoAttention 4.0 0.0 0.0 2.0 0.0 0.0 2.0 2.0 1.2 1.6 3.3 28.0 24.0 5.2
SnapKV 94.0 28.0 0.0 6.0 2.0 0.0 0.0 0.0 41.6 11.4 46.0 58.0 48.0 25.8
Quest 52.0 56.0 0.0 28.0 0.0 0.0 4.5 7.0 30.8 12.4 46.0 6.0 18.0 20.1
SparQ 4.0 0.0 0.0 2.0 0.0 0.0 0.5 0.0 20.0 30.4 35.3 46.0 40.0 13.7
Hybrid 98.0 86.0 72.0 78.0 92.0 30.0 80.0 27.5 89.2 31.8 71.3 72.0 56.0 68.0
RocketKV 96.0 88.0 84.0 76.0 96.0 10.0 81.5 53.5 89.2 33.6 70.7 66.0 54.0 69.1
512 Exact-TopK 100.0 98.0 96.0 96.0 100.0 82.0 92.5 64.0 91.2 36.6 78.0 64.0 58.0 81.3
DuoAttention 8.0 12.0 2.0 4.0 4.0 0.0 8.5 3.0 0.0 0.4 52.7 30.0 26.0 11.6
SnapKV 100.0 84.0 0.0 64.0 14.0 0.0 17.5 4.5 89.6 20.2 51.3 64.0 50.0 43.0
Quest 98.0 74.0 2.0 36.0 68.0 0.0 45.5 21.0 83.6 30.4 71.3 38.0 40.0 46.8
SparQ 98.0 94.0 56.0 76.0 92.0 16.0 39.0 23.0 78.0 38.4 70.0 72.0 56.0 62.2
Hybrid 100.0 96.0 84.0 94.0 100.0 74.0 85.0 51.5 88.8 37.0 71.3 68.0 58.0 77.5
RocketKV 100.0 100.0 96.0 96.0 100.0 36.0 95.0 67.5 90.0 33.6 72.0 68.0 60.0 78.0
1024 Exact-TopK 100.0 100.0 96.0 98.0 100.0 80.0 97.0 68.0 93.2 34.4 82.0 62.0 56.0 82.0
DuoAttention 12.0 12.0 0.0 4.0 4.0 0.0 8.5 5.0 1.2 0.8 62.0 32.0 30.0 13.2
SnapKV 100.0 88.0 6.0 92.0 34.0 2.0 63.5 21.5 91.2 29.8 60.0 68.0 58.0 54.9
Quest 100.0 90.0 30.0 80.0 78.0 24.0 80.0 49.5 92.4 32.8 73.3 66.0 54.0 65.4
SparQ 100.0 98.0 94.0 98.0 100.0 66.0 93.0 67.0 91.2 35.2 79.3 66.0 56.0 80.3
Hybrid 100.0 98.0 96.0 96.0 100.0 86.0 91.0 62.0 92.4 37.0 74.7 66.0 58.0 81.3
RocketKV 100.0 98.0 96.0 98.0 100.0 52.0 92.5 68.5 92.4 32.8 78.7 68.0 62.0 79.9
2048 Exact-TopK 100.0 100.0 96.0 98.0 100.0 80.0 96.5 69.0 92.8 34.4 82.7 66.0 58.0 82.6
DuoAttention 48.0 76.0 18.0 78.0 66.0 26.0 61.5 28.5 14.8 1.6 76.7 60.0 48.0 46.4
SnapKV 100.0 98.0 62.0 96.0 52.0 8.0 77.5 52.0 92.8 34.2 64.7 64.0 58.0 66.1
Quest 100.0 98.0 96.0 98.0 100.0 84.0 96.0 70.0 92.8 34.2 74.0 68.0 58.0 82.2
SparQ 100.0 100.0 96.0 98.0 100.0 80.0 97.0 68.5 92.8 34.0 82.0 66.0 58.0 82.5
Hybrid 100.0 100.0 96.0 98.0 100.0 80.0 98.0 67.0 91.2 36.6 79.3 62.0 58.0 82.0
RocketKV 100.0 100.0 96.0 98.0 100.0 74.0 95.5 68.0 92.8 33.2 80.7 64.0 60.0 81.7
4096 Exact-TopK 100.0 100.0 96.0 98.0 100.0 74.0 96.5 70.0 93.6 33.8 82.7 70.0 58.0 82.5
DuoAttention 98.0 100.0 88.0 96.0 100.0 64.0 97.0 74.5 85.6 35.0 82.0 66.0 58.0 80.3
SnapKV 100.0 98.0 94.0 98.0 88.0 26.0 92.5 65.0 94.0 34.4 74.0 66.0 60.0 76.1
Quest 100.0 100.0 96.0 98.0 100.0 78.0 96.5 69.5 93.6 34.0 80.0 66.0 58.0 82.3
SparQ 100.0 100.0 96.0 98.0 100.0 76.0 96.5 69.5 93.2 33.8 82.0 70.0 58.0 82.5
Hybrid 100.0 100.0 96.0 98.0 100.0 80.0 96.5 69.5 92.8 34.4 82.7 66.0 58.0 82.6
RocketKV 100.0 100.0 96.0 98.0 100.0 74.0 96.5 71.5 92.8 33.6 82.0 66.0 58.0 82.2
Table 17: Results of RULER, SeqLen=16000 for LongChat-7B-v1.5
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 100.0 96.0 94.0 98.0 42.0 84.5 50.0 82.4 18.0 72.0 64.0 44.0 72.7
256 Exact-TopK 100.0 98.0 98.0 86.0 100.0 54.0 74.0 44.5 75.6 38.4 70.0 58.0 44.0 72.3
DuoAttention 2.0 0.0 0.0 2.0 0.0 0.0 2.5 3.5 0.8 0.4 0.7 26.0 30.0 5.2
SnapKV 92.0 22.0 0.0 8.0 0.0 0.0 0.0 0.5 36.4 1.2 44.0 50.0 38.0 22.5
Quest 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.2 12.0 0.0 6.0 1.4
SparQ 6.0 2.0 0.0 0.0 0.0 0.0 0.0 0.0 2.8 0.0 12.0 2.0 14.0 3.0
Hybrid 96.0 86.0 24.0 38.0 22.0 0.0 36.5 11.5 76.8 15.2 39.3 50.0 40.0 41.2
RocketKV 100.0 98.0 96.0 86.0 96.0 38.0 60.0 34.5 35.6 12.8 64.0 54.0 46.0 63.1
512 Exact-TopK 100.0 100.0 98.0 84.0 100.0 54.0 78.5 50.0 76.8 29.8 73.3 58.0 44.0 72.8
DuoAttention 2.0 2.0 2.0 2.0 2.0 0.0 5.5 4.5 0.4 0.6 68.0 20.0 30.0 10.7
SnapKV 100.0 78.0 0.0 44.0 2.0 0.0 6.5 4.5 76.0 2.4 46.7 54.0 42.0 35.1
Quest 58.0 46.0 0.0 20.0 2.0 0.0 16.0 0.5 61.2 8.4 48.0 4.0 16.0 21.5
SparQ 46.0 54.0 0.0 6.0 10.0 0.0 5.5 0.5 64.8 32.6 22.0 30.0 36.0 23.6
Hybrid 100.0 96.0 90.0 68.0 62.0 8.0 72.0 47.5 78.8 19.6 52.0 60.0 46.0 61.5
RocketKV 100.0 100.0 96.0 94.0 98.0 44.0 77.0 58.0 64.0 11.2 70.0 64.0 46.0 70.9
1024 Exact-TopK 100.0 100.0 96.0 90.0 100.0 52.0 82.0 49.5 77.6 26.4 69.3 64.0 44.0 73.1
DuoAttention 6.0 8.0 2.0 2.0 2.0 0.0 7.0 4.5 1.2 0.6 65.3 28.0 34.0 12.4
SnapKV 100.0 86.0 0.0 76.0 14.0 0.0 25.5 11.0 77.2 9.8 57.3 56.0 46.0 43.0
Quest 94.0 98.0 8.0 74.0 42.0 0.0 55.5 19.0 80.0 26.0 58.7 36.0 34.0 48.1
SparQ 100.0 96.0 88.0 86.0 90.0 28.0 66.0 42.0 79.2 27.6 68.0 58.0 38.0 66.7
Hybrid 100.0 98.0 96.0 84.0 88.0 44.0 76.0 46.5 77.6 29.2 60.0 56.0 46.0 69.3
RocketKV 100.0 100.0 96.0 94.0 98.0 46.0 81.5 55.5 77.6 12.4 71.3 60.0 46.0 72.2
2048 Exact-TopK 100.0 100.0 96.0 92.0 98.0 50.0 82.5 51.0 81.6 25.0 71.3 62.0 44.0 73.3
DuoAttention 6.0 14.0 2.0 4.0 2.0 0.0 8.5 5.0 1.2 0.4 70.0 38.0 28.0 13.8
SnapKV 100.0 90.0 12.0 84.0 28.0 0.0 44.0 19.5 78.0 15.8 64.0 60.0 44.0 49.2
Quest 100.0 92.0 64.0 82.0 80.0 16.0 80.5 48.0 77.6 26.8 69.3 52.0 48.0 64.3
SparQ 100.0 100.0 94.0 90.0 100.0 54.0 82.0 53.5 77.6 21.2 74.7 64.0 44.0 73.5
Hybrid 100.0 100.0 98.0 88.0 98.0 56.0 81.5 52.0 78.8 29.6 69.3 62.0 46.0 73.8
RocketKV 100.0 100.0 98.0 92.0 98.0 44.0 82.5 58.5 78.8 15.0 72.0 62.0 48.0 73.0
4096 Exact-TopK 100.0 100.0 96.0 92.0 98.0 44.0 84.5 52.5 82.0 20.6 70.7 62.0 44.0 72.8
DuoAttention 82.0 90.0 38.0 90.0 58.0 4.0 68.5 46.0 34.8 0.4 70.7 58.0 46.0 52.8
SnapKV 100.0 96.0 46.0 92.0 42.0 4.0 67.5 37.0 78.4 20.8 65.3 60.0 44.0 57.9
Quest 100.0 100.0 98.0 92.0 96.0 62.0 81.5 53.5 78.4 26.6 68.7 62.0 44.0 74.1
SparQ 100.0 100.0 96.0 92.0 98.0 48.0 81.5 52.0 81.6 22.4 72.0 62.0 44.0 73.0
Hybrid 100.0 100.0 96.0 92.0 98.0 52.0 83.0 53.0 79.2 26.4 70.7 62.0 44.0 73.6
RocketKV 100.0 100.0 96.0 92.0 96.0 40.0 82.0 51.5 79.2 18.8 73.3 62.0 42.0 71.8
Table 18: Results of RULER, SeqLen=24000 for LongChat-7B-v1.5
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 100.0 98.0 90.0 84.0 18.0 60.5 38.5 79.6 5.8 70.0 50.0 36.0 63.9
256 Exact-TopK 100.0 98.0 98.0 88.0 78.0 16.0 46.5 39.0 60.0 24.8 64.7 46.0 38.0 61.3
DuoAttention 0.0 0.0 2.0 4.0 0.0 0.0 2.0 3.0 1.2 0.4 0.0 20.0 24.0 4.4
SnapKV 90.0 10.0 0.0 6.0 0.0 0.0 0.0 0.0 44.4 1.0 26.0 42.0 26.0 18.9
Quest 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 16.0 0.0 2.0 1.4
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 2.0 0.2
Hybrid 78.0 40.0 0.0 22.0 0.0 0.0 5.0 1.0 64.4 7.8 47.3 28.0 28.0 24.7
RocketKV 100.0 100.0 94.0 88.0 64.0 6.0 46.0 29.5 63.2 2.6 56.7 50.0 38.0 56.8
512 Exact-TopK 100.0 100.0 98.0 88.0 78.0 22.0 49.5 38.5 63.6 19.6 68.7 46.0 36.0 62.1
DuoAttention 2.0 4.0 4.0 4.0 2.0 0.0 4.5 4.0 0.4 0.6 55.3 22.0 22.0 9.6
SnapKV 100.0 28.0 0.0 28.0 2.0 0.0 2.0 3.0 70.4 2.0 30.0 48.0 30.0 26.4
Quest 96.0 0.0 0.0 0.0 0.0 0.0 0.5 1.5 32.8 0.6 50.7 2.0 12.0 15.1
SparQ 10.0 0.0 0.0 2.0 0.0 0.0 0.0 0.0 6.4 12.2 34.0 20.0 18.0 7.9
Hybrid 90.0 86.0 70.0 64.0 44.0 0.0 34.0 28.0 64.0 17.6 51.3 50.0 42.0 49.3
RocketKV 100.0 100.0 92.0 86.0 68.0 6.0 55.5 41.0 70.8 2.8 59.3 50.0 32.0 58.7
1024 Exact-TopK 100.0 100.0 98.0 86.0 82.0 20.0 58.0 35.5 66.8 12.2 70.7 46.0 34.0 62.2
DuoAttention 4.0 4.0 4.0 6.0 2.0 0.0 4.5 4.5 0.4 0.6 58.7 16.0 24.0 9.9
SnapKV 100.0 56.0 0.0 52.0 4.0 0.0 9.0 7.5 71.2 3.2 35.3 50.0 28.0 32.0
Quest 98.0 44.0 2.0 72.0 6.0 0.0 40.0 17.0 73.6 13.2 56.0 20.0 34.0 36.6
SparQ 86.0 62.0 0.0 20.0 12.0 0.0 4.5 5.5 76.8 27.6 26.0 48.0 34.0 31.0
Hybrid 100.0 98.0 90.0 88.0 76.0 10.0 40.0 36.5 72.0 26.2 52.0 50.0 42.0 60.1
RocketKV 100.0 100.0 96.0 86.0 80.0 8.0 56.5 38.5 73.2 2.6 64.7 48.0 36.0 60.7
2048 Exact-TopK 100.0 100.0 98.0 88.0 84.0 22.0 60.0 36.5 66.8 9.4 69.3 48.0 34.0 62.8
DuoAttention 4.0 4.0 4.0 6.0 2.0 0.0 4.5 4.5 0.4 0.4 57.3 22.0 24.0 10.2
SnapKV 100.0 72.0 2.0 66.0 12.0 0.0 18.0 15.0 77.6 5.0 48.0 50.0 34.0 38.4
Quest 90.0 74.0 42.0 68.0 50.0 0.0 57.5 24.5 79.2 34.2 56.7 44.0 36.0 50.5
SparQ 100.0 100.0 98.0 90.0 82.0 10.0 57.0 38.5 63.6 11.0 65.3 48.0 38.0 61.6
Hybrid 100.0 100.0 92.0 86.0 78.0 28.0 53.0 39.5 76.8 21.4 62.7 46.0 36.0 63.0
RocketKV 100.0 100.0 98.0 92.0 88.0 18.0 59.0 44.5 76.8 2.2 66.7 50.0 34.0 63.8
4096 Exact-TopK 100.0 100.0 98.0 90.0 84.0 20.0 63.5 36.5 68.4 7.6 72.0 48.0 36.0 63.4
DuoAttention 16.0 14.0 4.0 10.0 2.0 0.0 6.5 5.0 0.4 0.6 65.3 28.0 22.0 13.4
SnapKV 100.0 86.0 18.0 74.0 24.0 2.0 35.0 21.0 78.0 8.4 56.7 48.0 34.0 45.0
Quest 100.0 98.0 96.0 90.0 78.0 10.0 65.0 48.0 80.0 25.2 64.0 50.0 36.0 64.6
SparQ 100.0 100.0 98.0 90.0 82.0 18.0 61.0 38.0 65.2 7.0 69.3 48.0 36.0 62.5
Hybrid 100.0 100.0 98.0 88.0 82.0 22.0 58.5 38.5 67.2 12.2 66.0 46.0 34.0 62.5
RocketKV 100.0 100.0 98.0 92.0 88.0 16.0 65.0 39.0 72.8 3.6 68.7 50.0 32.0 63.5
Table 19: Results of RULER, SeqLen=32000 for LongChat-7B-v1.5
Token Budget
Method
single1
single2
single3
multikey1
multikey2
multikey3
multivalue
multiquery
vt
cwe
fwe
qa-1
qa-2
avg
N/A Full-KV 100.0 94.0 80.0 78.0 68.0 4.0 58.5 31.0 47.6 38.6 70.7 22.0 34.0 55.9
256 Exact-TopK 100.0 88.0 82.0 78.0 70.0 8.0 49.0 25.0 39.2 53.4 57.3 18.0 30.0 53.7
DuoAttention 0.0 2.0 4.0 4.0 0.0 0.0 1.0 2.5 0.4 0.4 0.7 20.0 26.0 4.7
SnapKV 80.0 2.0 0.0 2.0 0.0 0.0 0.0 0.0 13.2 0.4 13.3 24.0 32.0 12.8
Quest 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 4.7 0.0 0.0 0.4
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Hybrid 60.0 4.0 0.0 8.0 0.0 0.0 1.0 0.0 38.0 2.0 14.0 8.0 22.0 12.1
RocketKV 100.0 78.0 52.0 66.0 50.0 2.0 34.5 16.5 34.4 5.2 45.3 20.0 28.0 40.9
512 Exact-TopK 100.0 90.0 84.0 78.0 72.0 8.0 51.0 27.5 43.2 52.4 65.3 20.0 30.0 55.5
DuoAttention 2.0 6.0 8.0 4.0 0.0 0.0 2.5 1.5 0.0 0.4 33.3 20.0 28.0 8.1
SnapKV 100.0 14.0 0.0 12.0 0.0 0.0 0.5 1.0 34.4 1.0 11.3 24.0 30.0 17.6
Quest 24.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.4 0.4 18.7 0.0 10.0 4.1
SparQ 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 6.8 1.8 10.0 2.0 14.0 2.7
Hybrid 76.0 28.0 2.0 34.0 0.0 0.0 13.5 0.5 35.2 14.0 34.7 20.0 32.0 22.3
RocketKV 98.0 84.0 78.0 74.0 66.0 2.0 45.0 29.0 37.6 6.0 52.0 24.0 32.0 48.3
1024 Exact-TopK 100.0 90.0 84.0 78.0 72.0 6.0 50.0 26.0 44.0 49.8 68.0 20.0 32.0 55.4
DuoAttention 2.0 6.0 8.0 4.0 2.0 0.0 2.5 2.0 0.4 0.4 33.3 20.0 24.0 8.0
SnapKV 100.0 28.0 0.0 30.0 4.0 0.0 4.5 4.0 38.4 4.0 14.7 24.0 34.0 22.0
Quest 96.0 20.0 0.0 8.0 0.0 0.0 9.5 1.5 35.6 3.4 42.0 2.0 18.0 18.2
SparQ 54.0 8.0 0.0 0.0 0.0 0.0 0.0 0.0 20.8 33.0 22.0 18.0 34.0 14.6
Hybrid 94.0 64.0 72.0 70.0 12.0 0.0 54.0 17.5 40.0 29.2 42.7 22.0 30.0 42.1
RocketKV 98.0 94.0 80.0 74.0 64.0 2.0 49.0 32.5 41.2 9.0 55.3 24.0 34.0 50.5
2048 Exact-TopK 100.0 90.0 82.0 78.0 68.0 6.0 51.5 27.0 47.2 47.6 73.3 20.0 30.0 55.4
DuoAttention 2.0 6.0 8.0 4.0 2.0 0.0 2.5 2.0 0.8 0.2 35.3 18.0 24.0 8.1
SnapKV 100.0 60.0 2.0 50.0 6.0 2.0 11.5 13.0 38.4 7.8 22.7 20.0 32.0 28.1
Quest 88.0 90.0 0.0 46.0 14.0 0.0 50.5 17.5 50.0 26.6 52.0 20.0 26.0 37.0
SparQ 96.0 58.0 62.0 78.0 34.0 0.0 34.0 20.5 41.2 46.2 60.0 22.0 28.0 44.6
Hybrid 100.0 86.0 86.0 76.0 44.0 4.0 52.0 30.5 40.4 46.8 58.0 20.0 28.0 51.7
RocketKV 100.0 94.0 86.0 74.0 72.0 2.0 56.0 34.5 41.6 13.6 68.7 20.0 34.0 53.6
4096 Exact-TopK 100.0 94.0 82.0 78.0 70.0 6.0 53.0 30.0 46.8 41.2 71.3 20.0 32.0 55.7
DuoAttention 6.0 8.0 8.0 6.0 2.0 2.0 2.5 2.5 1.6 0.0 49.3 18.0 26.0 10.1
SnapKV 100.0 76.0 6.0 70.0 8.0 0.0 21.0 17.5 40.8 12.4 37.3 20.0 32.0 33.9
Quest 100.0 94.0 66.0 78.0 54.0 0.0 58.0 27.0 45.2 44.8 57.3 22.0 26.0 51.7
SparQ 100.0 94.0 84.0 78.0 58.0 2.0 49.5 26.5 45.6 47.4 72.0 20.0 32.0 54.5
Hybrid 100.0 90.0 90.0 76.0 70.0 6.0 49.5 30.5 43.2 44.2 66.0 20.0 32.0 55.2
RocketKV 100.0 92.0 86.0 78.0 76.0 4.0 57.5 38.0 43.2 17.8 70.7 16.0 36.0 55.0